### v1

# When can isotropy help adapt LLMs’ next word prediction to numerical domains?

Abstract
Recent studies have shown that vector representations of contextual embeddings learned by pre-trained large language models (LLMs) are effective in various downstream tasks in numerical domains. Despite their significant benefits, the tendency of LLMs to hallucinate in such domains can have severe consequences in applications such as energy, nature, finance, healthcare, retail and transportation, among others. To guarantee prediction reliability and accuracy in numerical domains, it is necessary to open the black-box and provide performance guarantees through explanation. However, there is little theoretical understanding of when pre-trained language models help solve numeric downstream tasks. This paper seeks to bridge this gap by understanding when the next-word prediction capability of LLMs can be adapted to numerical domains through a novel analysis based on the concept of isotropy in the contextual embedding space. Specifically, we consider a log-linear model for LLMs in which numeric data can be predicted from its context through a network with softmax in the output layer of LLMs (i.e., language model head in self-attention). We demonstrate that, in order to achieve state-of-the-art performance in numerical domains, the hidden representations of the LLM embeddings must possess a structure that accounts for the shift-invariance of the softmax function. By formulating a gradient structure of self-attention in pre-trained models, we show how the isotropic property of LLM embeddings in contextual embedding space preserves the underlying structure of representations, thereby resolving the shift-invariance problem and providing a performance guarantee. Experiments show that different characteristics of numeric data and model architecture could have different impacts on isotropy.
1 Introduction
Large language models have demonstrated impressive success in adapting to various downstream tasks in numerical domains, such as finance Garza and Mergenthaler-Canseco (2023); Yu et al. (2023), energy Gao et al. (2024), climate science Jin et al. (2024), healthcare Wang and Zhang (2024), wireless communications Xu et al. (2024), synthetic tabular generation Dinh et al. (2022); Borisov et al. (2023); Xu et al. (2024), among others. Inspired by the success of pre-trained LLMs, several methods have been developed recently in Gruver et al. (2024); Dooley et al. (2023); Nie et al. (2023); Rasul et al. (2024); Woo et al. (2024); Jin et al. (2024); Ansari et al. (2024) by adapting LLM to numerical domains. For many of these numeric downstream tasks, training a linear classifier on top of the hidden-layer representations generated by the pre-trained LLMs have already shown near state-of-the-art performance Jin et al. (2024); Ansari et al. (2024). However, these models in Gruver et al. (2024); Dooley et al. (2023); Nie et al. (2023); Rasul et al. (2024); Woo et al. (2024); Jin et al. (2024); Ansari et al. (2024) are treated as ‘black-box’ where numeric forecasts are controlled by complex nonlinear interactions between many parameters. This makes it difficult to understand how models arrive at their predictions and makes it challenging for users to trust the model outputs.
LLMs’ tendency to hallucinate can have serious consequences in critical numeric applications. For example, prediction errors in fraud detection in finance can lead to huge financial losses and errors in protection onset of sepsis or cardiac arrest in healthcare can result in patient deaths. Thus, to guarantee prediction reliability and accuracy in numerical domains, it is necessary to open the black-box and provide performance guarantees through explaination. Although recent empirical studies Jin et al. (2024); Nie et al. (2023); Liu et al. (2024) demonstrate the benefits of vector representations of embedding learned by LLMs in various numeric downstream tasks, there is little theoretical understanding of their empirical success. Thus, a fundamental question arises: “when (or how) can the next-word prediction capability of LLMs be effectively adapted to numerical domains?"
The main contribution of this paper is to provide an approach to answer this question by exploiting the isotropic property of LLM hidden representations in the contextual embedding space. Isotropy refers to the geometric property wherein vector representations in the embedding space are uniformly distributed in all directions, a characteristic critical for maintaining the expressiveness of the embedding space Arora et al. (2016); Mu and Viswanath (2018). To achieve state-of-the-art performance in numerical domains, we show that the hidden representations of LLMs must exhibit a structured form in contextual embedding space that accounts for the shift-invariance of the softmax function (i.e., the softmax output remains unchanged when all logits are shifted by a constant). Without such structure, the model can shift the logits while keeping the training loss unchanged, thereby leaving the logits ineffective for numeric downstream tasks. By formulating a gradient structure of self-attention in pre-trained models, we show how the isotropic property of LLM embeddings in contextual embedding space preserves the underlying structure of representations, thereby resolving the shift-invariance problem of the softmax function.
Our key contributions include: (i) We consider a log-linear model for LLMs and demonstrate theoretically why hidden representations must exhibit structure to address the shift-invariance problem of the softmax function. (ii) We take a deeper look into the hidden representations of pre-trained models and show how isotropy preserves the structural integrity of representations. In particular, we derive an upper bound for the Jacobian matrix which collects all first-order partial derivatives of self-attention with respect to the input pattern and show that $m$ largest eigenvectors of the LLM hidden representations minimize the gradient norm of self-attention. Then by projecting the representations into lower dimensions using these $m$ largest eigenvectors, we find the isotropy within the clusters in the contextual embedding space. (iii) Finally, we provide a comprehensive evaluation across $12$ real and $10$ synthetic
Figure 1: Time series tokenization.
time series datasets over $6$ different LLMs.
2 Problem Setup in Numerical Domains
Time Series Tokens and Similarity Measure. Similar to next-word prediction by LLMs, the next-value prediction in the numerical domain can be modeled by time series forecasting techniques Jin et al. (2024); Ansari et al. (2024) which are widely adopted in the machine learning literature. Formally, given a time series $\textbf{x}_{1:T+L}=[x_{1},\ldots,x_{T},\ldots,x_{T+L}]$, where the first $T$ time instances give the historical context, the next $L$ time instances constitute the forecast region, and $x_{t}\in\mathbb{R}$ is the observation of each time instance, we are interested in predicting the joint distribution of next $L$ time instances, $p(\textbf{x}_{T+1:T+L}|\textbf{x}_{1:T})$. Since, the pre-trained models operate on tokens from a finite vocabulary, using them for time series data requires mapping the observations to a finite set of tokens. Based on different numeric applications and LLM architectures, various tokenization techniques, e.g., quantization and scaling Ansari et al. (2024); Rasul et al. (2024), patching Woo et al. (2024); Jin et al. (2024); Nie et al. (2023), and adaptation of language model tokenizer in numeric domians Gruver et al. (2024); Dooley et al. (2023), can be applied to tokenize the time series and create a time series vocabulary $\mathcal{V}$ of $N$ time series tokens, i.e., $|\mathcal{V}|=N$, as shown in Figure 1. Then, the realization of the next $L$ time instances can be obtained by autoregressively sampling from the predicted distribution $p(k_{T+l+1}\,|\,\mathbb{k}_{1:T+l})$, for $l\in\{1,\ldots,L\}$, where $\mathbb{k}_{1:T+l}$ is the tokenized time series and $k_{i}$ be a time series token in time series vocabulary $|\mathcal{V}|$.
Let $\tilde{\mathbb{\Psi}}(k_{i})=\{\mathbb{\psi}_{1}(k_{i}),\mathbb{\psi}_{2}(k_{i}),\ldots\}$ be the set of all LLM contextual embedding instances of time series token $k_{i}$. Here, different contexts in the time series sequences yield different LLM embeddings of $k_{i}$. By constructing $\sum_{k}|\tilde{\mathbb{\Psi}}(k)|=|\mathcal{V}|$, we define the inter-token cosine similarity as,
$\zeta_{\textrm{cos}}\triangleq\mathbb{E}_{i\neq j}[\cos(\mathbb{\psi}(k_{i}),\mathbb{\psi}(k_{j}))],\vspace{-0.9in}$ (1)
where $\mathbb{\psi}(k_{i})$ and $\mathbb{\psi}(k_{j})$ are random samples from $\tilde{\mathbb{\Psi}}(k_{i})$. The expectation is taken over all pairs of different tokens. The inter-token cosine similarity metric describes the similarity between different tokens based on the contexts. For ease of reading, we express $T+l$ as $T_{l}$ and $T+l+1$ as $T_{l+1}$ for the rest of the paper.
Model. We consider a general pre-trained model for numerical data and open the black box of the pre-trained model. Specifically, we assume that the observation probability of $k_{T_{l+1}}$ given $\mathbb{k}_{1:T_{l}}$ satisfies the log-linear model Arora et al. (2016)
$p^{*}(k_{T_{l+1}}\!=\!i\,\,|\,\,\mathbb{k}_{1:T_{l}})\propto\exp(\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle),$ (2)
where $\mathbb{\psi}^{*}(k_{i})\in\mathbb{R}^{D}$ is a vector that only depends on the time series token $k_{i\in\mathcal{V}}$, and $\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}})$ is a function that encodes the tokenized time series sequence $\mathbb{k}_{1:T_{l}}$ into a vector in $\mathbb{R}^{D}$. The log-linear modeling aligns with the commonly used LLMs networks whose last layer is typically a softmax layer. Moreover, we do not consider any prior distribution for input, which makes our model more general than previous latent models Arora et al. (2016); Wei et al. (2021).
To define the numeric downstream task, let $z^{*}_{i}(k,l)\!\!:=\!\!\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle$ be the $i$-th logit of the ground-truth model, and assume that the numeric downstream tasks are defined by a function of the logits, i.e., $f^{*}(\mathbb{z^{*}})$. Also let $Z^{*}(k,l)\!\!=\!\!\sum_{i=1}^{N}\exp(z^{*}_{i}(k,l))\!=\!\sum_{i=1}^{|\mathcal{V}|}\exp(\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle)$ be the partition function Arora et al. (2016), i.e., normalization factor. In LLMs, the partition function is often used to normalize the output probabilities of the model, ensuring that they sum to $1$. Then, the normalized ground-truth model $\forall i\!\in\!\mathcal{V}$ is given by
$p(k_{T_{l+1}}\!\!=i\,\,|\,\,\mathbb{k}_{1:T_{l}})$ $=\!\frac{\exp(\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle)}{\sum_{i=1}^{|\mathcal{V}|}\exp(\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle)}=\frac{\exp(z^{*}_{i}(k,l))}{Z^{*}(k,l)}$
Since we do not know the ground-truth model in reality, we do not have access to the ground-truth model $\mathbb{\psi}^{*}(k_{i})$ and $\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}})$. Instead, we only have access to the student model $\mathbb{\psi}(k_{i})$ and $\mathbb{\mathbb{\psi}}(\mathbb{k}_{1:T_{l}})$ that aims to achieve low pre-training loss. We can define the student logits as $\mathbb{z}(k,l)\!\!:=\!\!\{\langle\mathbb{\mathbb{\psi}}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}(k_{i})\rangle\}_{i=1}^{|\mathcal{V}|}$. Intuitively, $\mathbb{z}$ are the contextualized representations learned by the student-model during pre-training. Then, the solution of the downstream task is to learn a function $f(k,l)$. Then, the output of the student model $\forall i\!\in\!\mathcal{V}$ can be defines as
$p\,(k_{T_{l+1}}=i\,\,|\,\,\mathbb{k}_{1:T_{l}})\!$ $=\!\frac{\exp(\langle\mathbb{\mathbb{\psi}}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}(k_{i})\rangle)}{Z(k,l)}.\vspace{-1.5in}$ (3)
Loss Function. As typical in language models, we use the categorical distribution over the elements in the time series vocabulary $\mathcal{V}$ as the output distribution $p(k_{T_{l+1}}|\mathbb{k}_{1:T_{l}})$, for $l\in\{1,\ldots,L\}$, where $\mathbb{k}_{1:T_{l}}$ is the tokenized time series. The student model is trained to minimize the cross entropy between the distribution of the tokenized ground truth label and the predicted distribution. The loss function for a single sequence of tokenized time series is given by Ansari et al. (2024); Wu et al. (2023)
$\mathcal{L}$ $=\!-\!\sum_{l=1}^{L+1}\!\sum_{i=1}^{|\mathcal{V}|}p^{*}(k_{T_{l+1}}\!=\!i\!\,\,|\!\,\,\mathbb{k}_{1:T_{l}})\log p(k_{T_{l+1}}\!=\!j\,\,\!|\!\,\,\mathbb{k}_{1:T_{l}})$
$=\sum_{l=1}^{L+1}\!\sum_{i=1}^{|\mathcal{V}|}\!\mathcal{D}_{\textrm{KL}}(p^{*}(k_{T_{l+1}}\!=\!i|\mathbb{k}_{1:T_{l}})\|\,p(k_{T_{l+1}}\!=\!j|\mathbb{k}_{1:T_{l}}))+\!H(p^{*}(k_{T_{l+1}}\!=\!j|\mathbb{k}_{1:T_{l}})),\vspace{-5in}$ (4)
where $p(k_{T_{l+1}}\!\!=\!i\,|\,\mathbb{k}_{1:T_{l}})$ is the categorical distribution predicted by the student model parametrized by $v_{1:T_{l}}$, $p^{*}(k_{T_{l+1}}\!\!=i|\mathbb{k}_{1:T_{l}})$ is the distribution of ground-truth model, $\mathcal{D}_{\textrm{KL}}$ is the KL divergence, and $H(p^{*}(k_{T_{l+1}}\!\!=\!i|\,\mathbb{k}_{1:T_{l}}))$ is the entropy of distribution $p^{*}(k_{T_{l+1}}\!=\!i|\mathbb{k}_{1:T_{l}})$ which is a constant. We assume that student model achieves a small loss so that the KL-divergence term in (4) is also small.
Downstream Numerical Task. We consider a simple downstream task whose prediction on categorical distribution is linear in $\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}})$, that is, $f^{*}(k,l)=\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),u^{*}\rangle=\sum_{i=1}^{|\mathcal{V}|}a^{*}_{i}z^{*}_{i}(k,l),$ where $u^{*}=\sum_{i=1}^{|\mathcal{V}|}a^{*}_{i}\mathbb{\psi}^{*}(k_{i})\in\mathbb{R}^{D}$ and $a_{j}$ is the coefficient. This model is still not sufficient to provide a performance guarantee to generalize to downstream task in unseen scenarios. However, the log probability difference is proportional to the difference in the value of the perfect model (i.e., ground-truth) $f^{*}(k,l)$. This allows the student model to alter the signs of $f^{*}(k,l)$ without resulting in a large KL divergence Wu et al. (2023). Then, it is more reasonable to model the numeric downstream task as
$f^{*}(k,l)$ $=\sum_{i=1}^{|\mathcal{V}|}a^{*}_{i}\sigma(z^{*}_{i}(k,l)-b^{*}_{i})=\sum_{i=1}^{|\mathcal{V}|}a^{*}_{i}\sigma(\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle-b^{*}_{i}),\vspace{-1.5in}$
where $\sigma$ is the ReLU function and $b^{*}_{j}$ denotes the threshold for the logits. The numeric downstream task only considers the logits that are above the threshold, and thus ignores all the entries with very small probabilities.
3 The Role of Isotropy in Adapting LLMs to Numerical Data
As previously discussed in Section 2, we consider LLM networks whose last layer is usually a softmax layer and the numeric downstream task is determined by the function of the logits. The underlying relation between the logits and softmax function determines the performance of the numeric downstream tasks. However, the softmax function is shift-invariant, that is, the output of the softmax function remains unchanged when all logits are shifted by a constant. Since we do not have any control over the logit shift of the student model on unseen data, good performance during training does not necessarily provide any performance guarantee for the numeric downstream task on unseen scenarios. This can be formalized in the following theorem and the proof is provided in Appendix A.
Theorem 3.1.
Let the logits of the ground-truth model be bounded. Then for any $f^{*}(k,l)$, there exists a set of functions $\{\hat{z}_{i}(k,l)\}_{i=1}^{|\mathcal{V}|}$ such that for all $k$ and $T_{l+1}$, the predictive distribution of the student model $\hat{p}\,(k_{T_{l+1}}\,|\,\mathbb{k}_{1:T_{l}})$ matches that of ground-truth model $p^{*}(k_{T_{l+1}}\,|\,\mathbb{k}_{1:T_{l}})$ and $\hat{f}(k,l)=0$. In other words, there exists a student model with the same pre-training loss as the ground-truth model, but its logits are ineffective for the numeric downstream tasks.
Theorem 3.1 demonstrates that without any structure in the hidden representations of LLM embeddings, the student model is able to shift the logits for any sample while keeping the pre-training loss unchanged and leaving logits ineffective for the numeric downstream tasks. Consequently, a theoretical guarantee for the numeric downstream task performance needs structure in the LLM representations learned by the pre-trained model. One way to prevent the shift-invariance problem from influencing the performance of the numeric downstream tasks is to keep the partition function stable. Let $\mathbb{\Psi}\!=\!(\mathbb{\psi}_{1}(k),\ldots,\mathbb{\psi}_{|\mathcal{V}|}(k))^{\top}\!\in\!\mathbb{R}^{|\mathcal{V}|\times D}$ be the hidden representations of input time series sequence. Then the stability of the partition function can be assessed through the isotropy in the contextual embedding space Arora et al. (2016); Mu and Viswanath (2018) as follows
$I(\{\mathbb{\psi}(k)\})=\frac{\min_{\psi(\mathbb{k})\in\mathcal{C}}Z(k,l)}{\max_{\psi(\mathbb{k})\in\mathcal{C}}Z(k,l)},$ (5)
where $\mathcal{C}=\mathbb{\Psi}^{\top}\mathbb{\Psi}$ is the input correlation matrix of input pattern and $l={1,\ldots,L}$. From (5), we can see that when the partition function is constant (i.e., stable) for different samples, $I(\{\mathbb{\psi}(k)\})$ becomes close to $1$ which indicates that the contextual embedding space $\{\mathbb{\psi}(k)\}$ is more isotropicArora et al. (2016); Mu and Viswanath (2018). Note that in (3), the probability of a value in any time instance is the exponential of the corresponding logit $z_{i}(k,l)$ divided by the partition function $Z(k,l)$. If the partition function remains stable for different samples, the logits can be solely determined by the probabilities, thereby resolving the shift-invariance problem of the softmax function.
4 Study of isotropy in LLM hidden representations
Analysis settings. For illustrative purposes, in this section, we present two examples from a specific numerical domain, such as transport, and for a specific language model (i.e., GPT-2) that illustrate the conditions under which isotropy is preserved in LLM representations.
Figure 2: Visualization of transport Dataset 1 (Left) and Dataset 2 (Right) labels for two extreme cases of velocities, i.e., $10$ km/hour and $100$ km/hour.
In Section 5, we provide comprehensive evaluations across $22$ different datasets from various numerical domains and other $5$ different language models with different architectures and tokenization techniques. For this specific analysis, we select two datasets from transport as they are dynamic, noisy, time-varying, and thus hold all primary characteristics of the numerical data across various domains. Specifically, we use two ways of sharing signals which are typically used for signal transmission in transportation settings, as shown in Figure 2. We call these two datasets as “Dataset $1$” and “Dataset $2$”. The downstream task here is to predict the transport signalling property over wireless channel using LLM, where Dataset $1$ causes good downstream performance (i.e., near optimal signal prediction), while the Dataset $2$ causes bad downstream task performance (i.e., high error in signal prediction). We use NMSE as a performance metric for the numeric downstream task because it is widely used for signal prediction. We deploy the first six layers of GPT2 Radford et al. (2019) and use the datasets and simulation setups from Liu et al. (2024), which are the standard settings for time series forecasting. We predict $L=4$ future signaling properties based on the historical $T=16$ signal properties through time series forecasting using GPT2. The training and validation dataset contains $8,000$ and $1,000$ samples, respectively, with user velocities uniformly distributed between $10$ km/hour and $100$ km/hour. The test dataset contains ten velocities ranging from $10$ km/hour to $100$ km/hour, with $1,024$ samples for each velocity.
Table 1: The effective dimension $d(0.8)$
Layer 1 2 3 4 5 6
Dataset 1 4 4 4 4 4 4
Dataset 2 1 1 1 1 1 1
Effective Dimensions. In each layer of each model, we start with a data matrix $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times D}$, where $|\mathcal{V}|$ represents the number of tokens in the input time series sequence, and $D$ corresponds to the embedding dimension. We apply PCA to reduce the dimensionality from $D$ to $m$ i.e., $\mathbf{\tilde{A}}\in\mathbb{R}^{|\mathcal{V}|\times m}$. Then the fraction of variance captured by the reduced representation is given by: $r_{m}=\frac{\sum_{i=0}^{m-1}\sigma_{i}}{\sum_{i=0}^{D-1}\sigma_{i}}$ where $\sigma_{i}$ denotes the $i$-th largest eigenvalue of the covariance matrix of $\mathbf{A}$. We define the $\epsilon$-effective dimension as $d(\epsilon)\triangleq\arg\min_{m}r_{m}\geq\epsilon$. For instance, if $d(0.8)=3$, it means that three principal dimensions retain $80$% of the variance. A higher $d$ suggests a more isotropic space Cai et al. (2021), where information is spread across multiple dimensions rather than being concentrated in a narrow subspace. Table 1 presents the values of $d(0.8)$ for different layers and models. Surprisingly, GPT-2 has so few effective dimensions, with $d(0.8)=4$ for Dataset $1$ and $d(0.8)=4$ for Dataset $1$, for layers 1 through 6, as compared to its original embedding dimensions $D=768$. The reduced dimensionality suggests that GPT-2’s embeddings lie in a subspace defined by a very narrow cone Ethayarajh (2019), and hence, there would be hardly any model capacity. Surprisingly, these language models are as successful as they are in numerical domains, given that most of their embedding vectors are as similar to one another as they are. These observations motivate us to look deeper into the contextual embedding space.
4.1 Clusters in the Contextual Embedding Space
Let $G(\mathbb{\Psi})\!=\!(g_{1}(\mathbb{\Psi}),\ldots,g_{|\mathcal{V}|}(\mathbb{\Psi}))^{\top}\!:\mathbb{R}^{|\mathcal{V}|\times D}\mapsto\mathbb{R}^{|\mathcal{V}|\times D}$ be the function for self-attention, i.e., $g_{i}(\mathbb{\Psi})=\text{softmax}(\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\Psi}^{\top})\mathbb{\Psi}$, where $\mathbb{\Lambda}\!=\!\mathbb{W}_{Q}\mathbb{W}_{K}^{\top}\in\mathbb{R}^{D\times D}$, and $\mathbb{W}_{Q}\!\!\in\!\!\mathbb{R}^{D\times m}$, $\mathbb{W}_{K}\!\!\in\!\!\mathbb{R}^{D\times m}$ are the parameter matrices for the query and key matrices of self-attention. The lemma below contributes to understanding why the isotropic property of pre-trained LLMs help to generalize to numerical domains. The proof of this lemma closely aligns with the approach in Kim et al. (2021). The proof of this lemma follows the analysis in Kim et al. (2021) is provided in Appendix B for completeness.
Lemma 4.1.
Consider the Jacobian matrix $\mathbb{J}=\left[\frac{\partial g_{i}(\mathbb{\Psi})}{\partial\mathbb{\psi}_{j}}\right]_{i,j=1}^{|\mathcal{V}|}$, which represents the gradient of the self-attention mapping $G(\mathbb{\Psi})$ with respect to the input time series token embeddings. Then the spectral norm of $\mathbb{J}$ satisfies $\|\mathbb{J}\|_{2}\leq|\mathbb{\Lambda}|_{2}\sum_{i=1}^{|\mathcal{V}|}\left(p_{i,i}+\frac{1}{2}\right)\left|\mathbb{\psi}_{i}-\sum_{j=1}^{|\mathcal{V}|}p_{i,j}\mathbb{\psi}_{j}\right|^{2}+\Delta$, where the residual term $\Delta$ is given by $\Delta=|\mathbb{\Lambda}|_{2}\sum_{i\neq j}^{|\mathcal{V}|}p_{i,j}\left|\mathbb{\psi}_{j}-\sum_{q=1}^{|\mathcal{V}|}p_{i,q}\mathbb{\psi}_{q}\right|^{2}+\frac{|\mathbb{\Lambda}|_{2}}{2}\sum_{j=1}^{|\mathcal{V}|}|\mathbb{\psi}_{j}|^{2}$, and the attention weights $p_{i,j}$ are defined as $p_{i,j}=\frac{\exp(\mathbb{\psi}_{i}^{\top}\mathbb{\Lambda}\mathbb{\psi}_{j})}{\sum_{k=1}^{|\mathcal{V}|}\exp(\mathbb{\psi}_{i}^{\top}\mathbb{\Lambda}\mathbb{\psi}_{k})}$.
From Lemma 4.1, we can see that to minimize the norm of the gradient $\|\mathbb{J}\|_{2}$, we essentially need to make $\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\sum_{j=1}^{|\mathcal{V}|}p_{i,j}\mathbb{\psi}_{j}\right|^{2}$ small. When $\mathbb{\Lambda}$ is small and all the input time series token embeddings are centered at the origin, $\sum_{i=1}^{|\mathcal{V}|}\mathbb{\psi}_{i}\!=\!0$, we have $\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}p_{i,:}\right|^{2}\approx\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}\right|^{2}$ (see Appendix B).
The theorem below shows that $\mathbb{\Lambda}$ minimizes the objective $\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}\right|^{2}$ and contains the $m$ largest eigenvectors of correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$ of time series token embeddings, where $m$ is the rank of $\mathbb{\Lambda}$. The proof of Theorem 4.1 is provided in Appendix C.
Theorem 4.1.
Let the eigenvalues of the correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$ be ordered as $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{D}$, and let $\gamma_{i}\in\mathbb{R}^{D}$ for $i=1,\ldots,D$ denote their associated eigenvectors. Then, the matrix $\mathbb{\Lambda}^{*}$ that minimizes the quantity $\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}\right|^{2}$ has the optimal form $\mathbb{\Lambda}=\sum_{i=1}^{m}\frac{1}{\lambda_{i}}\gamma_{i}\gamma_{i}^{\top}.$
Theorem 4.1 shows that the self-attention learns to perform a function for the numeric downstream tasks through training, which are closely related to the $m$ largest eigenvectors of LLM hidden representations. In other words, the self-attention mechanism effectively projects input time series tokens onto a low-dimensional contextual embedding space defined by the top eigenvectors of correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$.
Figure 3: Variations in GPT-2’s hidden representation for different datasets from the same domain: (a) PCA plot of contextual embedding space for transport Dataset 1. (b) PCA plot of contextual embedding space for transport Dataset 2.
Motivated by the findings from Lemma 4.1 and Theorem 4.1, we project the models’ hidden representations of into a lower-dimensional space by utilizing the $m\!=\!3$ largest eigenvectors through PCA, as shown in Figure 3. The three axes of the figure represent the first three principal components of the covariance matrix of LLM representations of each layer. For instance, in layer $6$, the first three principal components account for $76$% and $92$% of the total variance for Dataset $1$ and Dataset $2$, respectively. From Figure 3 a and Figure 3 b, we can see that there are disconnected or slightly overlapped islands that are far away from each other through layer 3 to layer 6. Note that the first principal dimension value spans from 00 to $2000$, significantly wider than the other $2$ dimensions, and dominates the total variance. A similar analogy can also be observed for transport Dataset 2 in Figure 3 b. In (1), the space isotropy is measured on pairs of arbitrary time series token representations, which could reside in two disconnected clusters. However, given that the variance is dominated by distances between clusters, such estimation would be biased by the inter-cluster distances. Hence, it is more meaningful to consider a per-cluster (i.e., local) investigation rather than a global estimate.
We start by performing clustering on the LLM representations in the contextual embedding space. There are various methods for performing clustering, such as $k$-means, DBSCAN Ester et al. (1996). We select $K$-means clustering method because it is reasonably fast in high embedding dimensions. We use the classical silhouette score analysis Rousseeuw (1987) to determine the number of clusters $|C|$ in the contextual embedding space (see Appendix D for details). Since each LLM contextual embedding instance $\mathbb{\psi}(k_{i})$ belongs to a particular cluster through clustering, the cosine similarity should be measured after shifting the mean to the origin Mu and Viswanath (2018). Accordingly, we subtract the mean for each cluster (i.e., centroid) and calculate the adjusted $\zeta_{\text{cos}}$ in Section 2. Assuming we have a total of $|C|$ clusters, let $\mathbb{\psi}_{c}(k_{i})=\{\mathbb{\psi}_{c}^{1}(k_{i}),\mathbb{\psi}_{c}^{2}(k_{i}),\dots\}$ be the set of token $k$’s contextual embeddings in cluster $c\in C$, and $\mathbb{\psi}_{c}(k_{i})$ be one random sample in $\mathbb{\psi}_{c}(k_{i})$. We define the adjusted inter-token cosine similarity as
$\zeta^{\prime}_{\textrm{cos}}\triangleq\mathbb{E}_{c}\left[\mathbb{E}_{i\neq j}\left[\cos\left(\bar{\mathbb{\psi}}_{c}(k_{i}),\bar{\mathbb{\psi}}_{c}(k_{j})\right)\right]\right],\vspace{-0.18in}$ (6)
where $\bar{\mathbb{\psi}}_{c}(k_{i})=\mathbb{\psi}_{c}(k_{i})-\mathbb{E}_{\mathbb{\psi}_{c}}[\mathbb{\psi}_{c}(k_{i})]$. Here $\mathbb{E}_{c}$ is the average over different clusters, and $\bar{\mathbb{\psi}}_{c}(k_{i})$ is the original contextual embedding shifted by the mean, with the mean taken over the samples in cluster $c$ Kim et al. (2021). The inter-token cosine similarity takes values between $-1$ and $1$. A value close to 0 indicates strong isotropy and ensures the existence of structure in the LLM representations.
Isotropy in Dataset 1. In this section, we provide an example of Dataset $1$ where a high prediction accuracy (i.e., low NMSE) is achieved by GPT2 based model. For instance, in Figure 4a, we compare the NMSE performance of our GPT2 based LLM with non-language models for different user velocities. From Figure 4 a, we can observe that the NMSE performance of all models gradually increased along with the increase in user velocity. This is because, with the increase in velocity, the transport signal characteristics rapidly change within a very short coherence time, resulting in increased prediction difficulty for the prediction model. The GPT2 based model consistently outperforms other baselines and demonstrates its high prediction accuracy.
Figure 4: (a) LLM outperforms all other baselines for all of the ten different velocities for Dataset 1. (b) Inter-type cosine similarities for Dataset $1$ with different velocities. $\zeta\textquoteright_{\textrm{cos}}$ are close to zero for all the layers, including layer 6, indicating that nearly perfect isotropy exists in the LLM embedding space for the Dataset $1$, which preserves the structure in the LLM’s hidden representations and causes the high prediction accuracy.
For illustrative purposes, we pick three user velocities: $10$ km/hour, $50$ km/hour, and $100$ km/hour, for isotropy assessment of Dataset $1$. The GPT2 based model achieves good NMSE performance for all of these three velocities, as shown in Figure 4a. We apply inter-type cosine similarity $\zeta^{\prime}_{\textrm{cos}}$ in (6) to measure the isotropy in GPT2 embedding space. From Figure 4b, we can see that the GPT2 has consistent near-zero cosine similarity values for all layers, including layer 6. This indicates that nearly perfect isotropy exists in the GPT2 embedding space for the Dataset $1$, which preserves the structure in the GPT2’s hidden representations and causes good downstream task performance.
Figure 5: (a) The NMSE performance of the LLM based model for Dataset 2 deteriorates significantly compared to Dataset 1. (b) Inter-type cosine similarities for Dataset $2$ for different velocities. Higher $\zeta_{\textrm{cos}}$ values indicate a weak isotropy (i.e., anisotropy) exists in the LLM embedding space which causes a lack of structure in the hidden representations, yielding poor prediction accuracy.
Isotropy in Dataset 2. In this section, we provide an example of Dataset $2$ where a poor prediction accuracy (i.e., high NMSE) is achieved by GPT2 based model. As shown in Figure 5 a, the NMSE performance fluctuates randomly for different velocities, while the NMSE performance for Dataset $1$ is gradually increasing with increase in the velocities. The NMSE performance for Dataset $2$ deteriorates significantly compared to the Dataset $1$. As before, with the three user velocities, the NMSE performance for Dataset $2$ for all of these velocities is worse as compared to Dataset $1$, as shown in Figure 5 a. From Figure 5 b, we can observe a weak isotropy (i.e., anisotropy) in the LLM embedding space for Dataset $2$, causing a lack of structure in the GPT2 hidden representations, and thus leading to bad downstream performance.
5 Experiments
Baselines. We consider popular pre-trained LLMs as the baselines for numeric downstream tasks, including PatchTST Nie et al. (2023), Lag-Llama Rasul et al. (2024), Moirai-1.0-R Woo et al. (2024), Chronos-T5 Ansari et al. (2024) and Choronos-Bolt (https://huggingface.co/autogluon/chronos-bolt-base). The models use different architectures, time series tokenization techniques and hyperparameters for numeric downstream tasks as summarized in Table 2.
Table 2: LLM models architectures, time series tokenization techniques and hyperparameter choices. $L$ stands for context length, $d_{h}$ for hidden layer dimension, $n_{L}$ for number of layers, $n_{H}$ for number of heads, and $\eta$ for learning rate.
Model Architecture Tokenization Technique Hyperparameters
Chronos-T5 Encoder-Decoder with autoregressive forecasting Scaling & Quantization Default
Chronos-Bolt Encoder-Decoder with multi-step forecasting Scaling & Quantization Default
PatchTST Vanilla Encoder Patching Patch length: 16, Stride: 8, $d_{h}=32,n_{L}=2,n_{H}=4$
Moirai Encoder Patching $L=1024$, Patch length: selected by dataset-specific validation
Lag-Llama Decoder Lag Feature $L=32$
Table 3: Real and Synthetic Datasets
Data Subset Domain Dataset 1 Dataset 2
Real Datasets Energy Australian Electricity – Queensland State Australian Electricity – South Australia
Weather Solar Radiation Rainfall
Finance Exchange Rate NN5 Weekly Cash Withdrawals
Healthcare Hospital Patient Counts COVID-19 Deaths
Transportation Transportation Signaling 1 Transportation Signaling 2
Retail Car Sales Dominick
Synthetic Datasets Linear DotProduct kernel (C=0) DotProduct kernel (C=1)
Seasonality Seasonality kernel (period = 0.5W) Seasonality kernel (period = 0.25H)
Trend RationalQuadratic kernel ($\alpha=1$) RationalQuadratic kernel ($\alpha=10$)
Non-Linear RBF kernel (length scale $=0.1$) RBF kernel (length scale $=1$)
Stochastic WhiteKernel (noise level = 0.1) WhiteKernel (noise level = 1)
Datasets. We conduct a comprehensive evaluation using $12$ different real time series datasets from various numerical domains, including energy, nature, finance, healthcare, retail and transportation (data sources can be found in Table 4 of Appendix E). We also illustrate our findings using KernelSynth Ansari et al. (2024) (see Algorithm 1 in Appendix E for details), a method that generates $10$ more synthetic datasets via Gaussian processes in Section 5. In Section 4.1, we show how the LLM hidden representation varies for two different datasets from the same numerical domain and its impact on downstream task performance. So we select two different datasets from each numerical domain (for good and bad downstream task performance examples) for qualitative and quantitative analysis as shown in Table 3.
5.1 Qualitative Analysis
In this section, we analyze the time series forecasting by the baseline LLMs qualitatively. We focus on synthetically generated time series for a controlled analysis of different types of time series patterns which belong to $5$ different domains, such as linear, seasonality, trend, non-linear and Gaussian perturbation.
Figure 6: NMSE vs isotropy analysis for $10$ different synthetic datasets of $5$ different domains.
We are particularly interested in the isotropic measurement in the LLM’s last layer as it is related to the logits and probabilistic inference. So all isotropic measure provided in this section is based on the last layer of the baselines.
We begin by analyzing time series forecasting performance (i.e., NMSE) for different baselines and its relation with isotropy in Figure 6. For instance, in Figure 6 b, the inter-type cosine similarity in hidden embedding spaces of Chronos-T5 for seasonality (Dataset 1) is $-0.00007$ and for seasonality (Dataset 2) is $0.0047$. This indicates that stronger isotropy exists (i.e., inter-type cosine similarity value is close to 00) for Chronos-T5 for seasonality (Dataset 1) which preserves the structure in the Chronos-T5’s hidden representations and causes good downstream task performance. On the other hand, a weaker isotropy exists (i.e., inter-type cosine similarity value is far from 00) for Chronos-T5 in seasonality (Dataset 2), causing a lack of structure in Chronos-T5’s hidden representations, and thus leading to bad downstream task performance. Moreover, the NMSE for seasonality (Dataset 1) is lower for Chronos-T5, PatchTST and Lag-Llma, and higher for Chronos-Bolt and Morai, while the NMSE is higher for all the baselines except Morai for non-linear (Dataset 2). A similar analysis can also be observed for all other synthetic datasets and baselines in Figure 6. This shows that the same dataset from any particular domain may cause different forecasting performances for different baselines, as it generates different contextual embedding spaces for language models with different architectures and tokenization techniques (See Appendix F for full illustration).
Next, we examine the influence of isotropy on forecasting performance in two important scenarios: a) different input context lengths, and b) different levels of noises in the input data. The first scenario is important as it provides an analysis on selecting proper input context length rather than selecting the length through random trials and errors. The second scenario is important as it gives us ideas on how the level of noise in noisy data impacts performance, since the data in the real world is mostly noisy.
Isotropy in different input context lengths. We first analyze the impacts of isotropy under varying input context lengths. We select Choronos-T5 as an example model and non-linear (Dataset 1) as an example dataset. In Figure 11 in Appendix F, we show how the hidden representations of Chronos-T5 vary for two different input context lengths, such as $L=500$ and $L=500$, for non-linear (Dataset 1), which gives an indication of having
Figure 7: NMSE vs isotropy comparison across different input context lengths for synthetic datasets.
different isotropic measures for different context lengths.
In Figure 7, we compare the NMSE vs isotropy across two different input context lengths, $L=500$ and $L=100$, for different synthetic datasets. As can be seen from the figure, the isotropy values vary across different input context lengths and datasets. For instance, in seasonality (Dataset 1), we have (NMSE$=0.0000066$, cosine similarity$=-0.00076$) and (NMSE$=0.0793$, cosine similarity$=0.0011$) for $L=500$ and $L=100$, respectively. The decrease in isotropy significantly increases the NMSE for the input context length $L=100$. In contrast, in Linear (Dataset 2), the isotropy increases for the input context length $L=100$, which causes the decreases in NMSE for chornos-T5. In practice, the input context length is often selected randomly or through trial and error, which may cause higher forecasting errors for different datasets. Isotropy analysis enables us to understand how varying input context lengths influence the hidden representations of the language model. This insight helps guide improvements in forecasting performance by examining the isotropic properties of the contextual embedding space.
Isotropy in varying noise levels in datasets. Next, we focus on the second scenario to see the impact of noisy datasets on LLM’s performance. Again, we use the Chornos-T5 as an example language model. In Figure 8, we compare the NMSE vs isotropy across two different noise levels, one without noise, and the other with Gaussian noise with $\sigma=0.05$ standard deviation. From Figure 8, we can see consistently lower isotropy (i.e., inter-type cosine similarity far from 00) for all noisy synthetic datasets as compared to the datasets without noise. For instance, in trend (Dataset 2),
Figure 8: NMSE vs isotropy comparison across different noise levels in synthetic datasets.
we have (NMSE$=0.000024$, cosine similarity$=-0.00022$) and (NMSE$=0.0012$, cosine similarity$=0.0040$) for $\sigma=0$ and $\sigma=0.05$, respectively. The decrease in isotropy significantly increases the NMSE for the noisy dataset. In practice, the environments of many real numerical domains, such as nature and energy, are noisy and dynamic. In such an environment, it is not always possible to measure the noise in real time and take the necessary steps to clean the input time series for better performance. However, it is always possible to measure the isotropy from LLM hidden representations, which can be used as a measure of noise in the input datasets, and thus, help to improve the forecasting performance.
5.2 Quantitative Analysis
In Figure 9, we analyze the time series forecasting performance for 12 real datasets with different baselines and its relation with isotropy. For instance, in Figure 9 e, the inter-type cosine similarity in hidden embedding spaces of Morai for retail (Dataset 1) is $0.002$ and for retail (Dataset 2) is $0.1931$. This indicates that stronger isotropy exists for Morai for retail (Dataset 1) which preserves the structure in the Morai’s hidden representations and causes good downstream task performance. On the other hand, a weaker isotropy exists for Morai in retail (Dataset 2),
Figure 9: NMSE vs isotropy analysis for $12$ different real datasets of $6$ different domains.
causing a lack of structure in Morai’s hidden representations, and thus leading to bad downstream task performance. Moreover, the NMSE for retail (Dataset 2) is lower for all baselines except the Morai as compared to retail (Dataset 1), while the NMSE for energy (Dataset 1) is lower for all baselines except Lag-Llma as compared to energy (Dataset 2). A similar analysis can also be observed for all other real datasets and baselines in Figure 9. This again shows that a dataset may have different impacts on forecasting performances for different baselines, and isotropy can be used as a measure of how the dataset generates different contextual embedding spaces for different language models based on their architectures and tokenization techniques.
Finally, in Figure 10, we compare the NMSE vs isotropy for varying input context lengths to observe its impact on the real datasets. We select Lag-Llma as our example model. We compare the results for two different input context length: 1) the recommended input context length $L=144$ and the reduced input context length $L=96$. As can be seen from the figure, the violation in recommended input context length
Figure 10: NMSE vs isotropy comparison across different input context lengths for real datasets.
by reducing it from $L=144$ to $L=96$ not only decreases the NMSE performances, but also increases for some datasets. For instance, the inter-type cosine similarity values become close to 00, i.e., from $0.1091$ to $0.0012$ and from $0.2014$ to $0.0396$, respectively, for nature (Dataset 2) and finance (Dataset 1), which in turn improves the NMSE performances.
6 Conclusion and Limitations
Isotropy in embeddings as studied here can serve as a foundation for future research on the deeper understanding of LLMs and their applications in various domains. Beyond isotropy, there could be other methods to approximate the partition function with a constant and make the logits useful for the numeric downstream tasks. Moreover, our isotropy study only ensured the existence of structure in the LLM hidden representations and provides a performance guarantee when the structure is preserved by isotropy. Improving the numeric downstream task performance when structure is not preserved in the LLM representations is a topic of future work.
Appendix A Proof of Theorem 3.1
Theorem A.1.
Let the logits of the ground-truth model be bounded. Then for any $f^{*}(k,l)$, there exists a set of functions $\{\hat{z}_{i}(k,l)\}_{i=1}^{|\mathcal{V}|}$ such that for all $k$ and $T_{l+1}$, the predictive distribution of the student model $\hat{p}\,(k_{T_{l+1}}\,|\,\mathbb{k}_{1:T_{l}})$ matches that of ground-truth model $p^{*}(k_{T_{l+1}}\,|\,\mathbb{k}_{1:T_{l}})$ and $\hat{f}(k,l)=0$. In other words, there exists a student model with the same pre-training loss as the ground-truth model, but its logits are ineffective for the numeric downstream tasks.
Proof.
We select $\tau\in\mathbb{R}$ such that $\forall k,T_{l+1}$, $\tau<\min_{j\in\mathcal{V}}b_{j}^{*}-\max_{j\in\mathcal{V}}z_{i}^{*}(k,l)$, and $\forall k,T_{l+1},\forall j\in\mathcal{V}$. By setting $\hat{z}_{j}(k,l):=z_{i}^{*}(k,l)+\tau$, we get $\forall j\in\mathcal{V},$
$\hat{z}_{j}(k,l)-b_{j}^{*}<z_{i}^{*}(k,l)+\min_{j\in\mathcal{V}}b_{j}^{*}-\max_{j\in\mathcal{V}}z_{i}^{*}(k,T_{l+1})-b_{j}^{*}\leq 0,$
this implies that $\sigma(\hat{z}_{j}(k,l)-b_{j}^{*})=0$. Hence, $\forall k,T_{l+1}$ and we have $\hat{f}(k,l)=0$. ∎
Appendix B Proof of Lemma 4.1
Lemma B.1.
Consider the Jacobian matrix $\mathbb{J}=\left[\frac{\partial g_{i}(\mathbb{\Psi})}{\partial\mathbb{\psi}_{j}}\right]_{i,j=1}^{|\mathcal{V}|}$, which represents the gradient of the self-attention mapping $G(\mathbb{\Psi})$ with respect to the input time series token embeddings. Then the spectral norm of $\mathbb{J}$ satisfies $\|\mathbb{J}\|_{2}\leq|\mathbb{\Lambda}|_{2}\sum_{i=1}^{|\mathcal{V}|}\left(p_{i,i}+\frac{1}{2}\right)\left|\mathbb{\psi}_{i}-\sum_{j=1}^{|\mathcal{V}|}p_{i,j}\mathbb{\psi}_{j}\right|^{2}+\Delta$, where the residual term $\Delta$ is given by $\Delta=|\mathbb{\Lambda}|_{2}\sum_{i\neq j}^{|\mathcal{V}|}p_{i,j}\left|\mathbb{\psi}_{j}-\sum_{q=1}^{|\mathcal{V}|}p_{i,q}\mathbb{\psi}_{q}\right|^{2}+\frac{|\mathbb{\Lambda}|_{2}}{2}\sum_{j=1}^{|\mathcal{V}|}|\mathbb{\psi}_{j}|^{2}$, and the attention weights $p_{i,j}$ are defined as $p_{i,j}=\frac{\exp(\mathbb{\psi}_{i}^{\top}\mathbb{\Lambda}\mathbb{\psi}_{j})}{\sum_{k=1}^{|\mathcal{V}|}\exp(\mathbb{\psi}_{i}^{\top}\mathbb{\Lambda}\mathbb{\psi}_{k})}$.
Proof.
According to the analysis, the gradient of $g_{i}(\boldsymbol{\Psi})$ with respect to the variable $\boldsymbol{\psi}_{j}$ is expressed as $J_{i,j}=\frac{\partial g_{i}(\mathbb{\Psi})}{\partial\mathbb{\psi}_{j}}=p_{i,j}I+\mathbb{\Psi}^{\top}Q^{i}\left(\mathbb{\Psi}\mathbb{\Lambda}\delta_{i,j}+E_{j,i}\mathbb{\Psi}\mathbb{\Lambda}^{\top}\right)$ where the matrix $Q^{i}$ is defined by $Q^{i}=\text{diag}(p_{i,:})-p_{i,:}p_{i,:}^{\top}$. Here, $p_{i,:}\in\mathbb{R}_{+}^{|\mathcal{V}|}$ corresponds to the $i$-th row of the probability matrix $\boldsymbol{P}$, $E_{j,i}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ denotes a matrix with a single entry at the $(j,i)$-th position and zeros elsewhere, and $\delta_{i,j}\in\{0,1\}$ is the Kronecker delta. We thus have
$\|\mathbb{J}\|_{2}$ $\leq\sum_{i,j=1}^{|\mathcal{V}|}|J_{i,j}|_{2}$
$\leq\sum_{i,j=1}^{|\mathcal{V}|}p_{i,j}+\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\Psi}^{\top}Q^{i}\mathbb{\Psi}|_{2}|\mathbb{\Lambda}|_{2}+\sum_{i,j=1}^{|\mathcal{V}|}|\mathbb{\Psi}^{\top}Q^{i}E_{j,i}\mathbb{\Psi}|_{2}|\mathbb{\Lambda}|_{2}$
$\leq|\mathcal{V}|+|\mathbb{\Lambda}|_{2}\sum_{i=1}^{|\mathcal{V}|}\left(\sum_{j=1}^{|\mathcal{V}|}p_{i,j}|\mathbb{\psi}_{j}|^{2}-\left|\sum_{j=1}^{|\mathcal{V}|}p_{i,j}\mathbb{\psi}_{j}\right|^{2}\right)+|\mathbb{\Lambda}|_{2}\sum_{i,j=1}^{|\mathcal{V}|}|\mathbb{\Psi}^{\top}Q^{i}e_{j}\mathbb{\psi}_{i}^{\top}|$
$\leq|\mathcal{V}|+|\mathbb{\Lambda}|_{2}\sum_{i=1}^{|\mathcal{V}|}\sum_{j=1}^{|\mathcal{V}|}p_{i,j}|\mathbb{\psi}_{j}-\sum_{q=1}^{|\mathcal{V}|}p_{i,q}\mathbb{\psi}_{q}|^{2}+|\mathbb{\Lambda}|_{2}\sum_{i,j=1}^{|\mathcal{V}|}p_{i,j}|\mathbb{\psi}_{i}^{\top}(\mathbb{\psi}_{j}-\mathbb{\Psi}^{\top}p_{i,:})|$
$\leq|\mathbb{\Lambda}|_{2}\sum_{i=1}^{|\mathcal{V}|}\left(p_{i,i}+\frac{1}{2}\right)|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}p_{i,:}|^{2}+|\mathcal{V}|+|\mathbb{\Lambda}|_{2}\sum_{i\neq j}^{|\mathcal{V}|}p_{i,j}|\mathbb{\psi}_{j}-\mathbb{\Psi}^{\top}p_{i,:}|^{2}+\frac{|\mathbb{\Lambda}|_{2}}{2}\sum_{j=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}|^{2}$
$=|\mathbb{\Lambda}|_{2}\sum_{i=1}^{|\mathcal{V}|}\left(p_{i,i}+\frac{1}{2}\right)|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}p_{i,:}|^{2}+|\mathcal{V}|+|\mathbb{\Lambda}|_{2}\sum_{i\neq j}^{|\mathcal{V}|}p_{i,j}\left|\mathbb{\psi}_{j}-\sum_{q=1}^{|\mathcal{V}|}p_{i,q}\mathbb{\psi}_{q}\right|^{2}+\frac{|\mathbb{\Lambda}|_{2}}{2}\sum_{j=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}|^{2}$
$=|\mathbb{\Lambda}|_{2}\sum_{i=1}^{|\mathcal{V}|}\left(p_{i,i}+\frac{1}{2}\right)|x_{i}-X^{\top}p_{i,:}|^{2}+|\mathcal{V}|+\Delta,$
where $\Delta=|\mathbb{\Lambda}|_{2}\sum_{i\neq j}^{|\mathcal{V}|}p_{i,j}\left|\mathbb{\psi}_{j}-\sum_{q=1}^{|\mathcal{V}|}p_{i,q}\mathbb{\psi}_{q}\right|^{2}+\frac{|\mathbb{\Lambda}|_{2}}{2}\sum_{j=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}|^{2}.$
The theorem below shows that $\mathbb{\Lambda}$ minimizing the objective $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}|^{2}$ contains the largest $m$ eigenvectors of the correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$ of input time series token embeddings where $m$ is the rank of $\mathbb{\Lambda}$.
Lemma 1 implies that one of the key components in the Jacobian’s upper bound takes the form $|\mathbb{\psi}_{i}-\sum_{j=1}^{|\mathcal{V}|}p_{i,j}\mathbb{\psi}_{j}|^{2}$. Consequently, during optimization, it is natural to aim for a reduction in the gradient magnitude, which motivates minimizing the expression $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\sum_{j=1}^{|\mathcal{V}|}p_{i,j}\mathbb{\psi}_{j}|^{2}$. This leads to understand the choice of $\mathbb{W}^{Q}$ and $\mathbb{W}^{K}$ that minimize $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\sum_{j=1}^{|\mathcal{V}|}p_{i,j}\mathbb{\psi}_{j}|^{2}$, which is equivalent to solving the optimization problem $\min_{|\mathbb{\Lambda}|_{F}\leq\rho}\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\sum_{j=1}^{|\mathcal{V}|}p_{i,j}\mathbb{\psi}_{j}|^{2}$, where the scalar constraint $\rho$ regulates the size of $\mathbb{\Lambda}$.
To proceed, we consider the objective in the scenario where $\rho$ is small. In this case, we can approximate the attention weights by $p_{i,j}\approx\frac{1}{|\mathcal{V}|}+\frac{1}{|\mathcal{V}|}\mathbb{\psi}_{i}^{\top}\mathbb{\Lambda}\mathbb{\psi}_{j}$. Now, we define the average of embedding as $\bar{\mathbb{\psi}}=\mathbb{\Psi}^{\top}1/|\mathcal{V}|$. It then follows that $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}p_{i,:}|^{2}=\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\bar{\mathbb{\psi}}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}|^{2}$. Assuming all input time series patterns are zero-centered, i.e., $\bar{\mathbb{\psi}}=0$, we have $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}|^{2}=\text{tr}\left((I-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda})^{2}\mathbb{\Psi}^{\top}\mathbb{\Psi}\right)$. Theorem 4.1 establishes that the optimal $\mathbb{\Lambda}$ that minimizes $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}|^{2}$ is spanned by the top $m$ eigenvectors of $\mathbb{\Psi}^{\top}\mathbb{\Psi}$, where $m$ equals the rank of $\mathbb{\Lambda}$.
Appendix C Proof of Theorem 4.1
Theorem C.1.
Let the eigenvalues of the correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$ be ordered as $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{D}$, and let $\gamma_{i}\in\mathbb{R}^{D}$ for $i=1,\ldots,D$ denote their associated eigenvectors. Then, the matrix $\mathbb{\Lambda}^{*}$ that minimizes the quantity $\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}\right|^{2}$ has the optimal form $\mathbb{\Lambda}=\sum_{i=1}^{m}\frac{1}{\lambda_{i}}\gamma_{i}\gamma_{i}^{\top}.$
Proof.
Given that $\mathbb{W}_{Q}\in\mathbb{R}^{D\times m}$ and $\mathbb{W}_{K}\in\mathbb{R}^{D\times m}$, it follows that the matrix $\mathbb{\Lambda}$ has rank $m$. Hence, we know $\min_{\mathbb{\Lambda}}\sum_{i=1}^{|\mathcal{V}|}\|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}\|^{2}\geq\sum_{q=m+1}^{|\mathcal{V}|}\lambda_{q}.$ Now, if we set $\mathbb{\Lambda}$ to $\mathbb{\Lambda}=\sum_{i=1}^{m}\frac{1}{\lambda_{i}}\gamma_{i}\gamma_{i}^{\top}$, then we obtain $\sum_{i=1}^{|\mathcal{V}|}\|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}\|^{2}=\text{tr}\left(\left(I-\sum_{i=1}^{m}\gamma_{i}\gamma_{i}^{\top}\right)^{2}\mathbb{\Psi}^{\top}\mathbb{\Psi}\right)=\sum_{q=m+1}^{D}\lambda_{q}.$
Therefore, the optimal solution $\mathbb{\Lambda}$ for minimizing $\sum_{i=1}^{|\mathcal{V}|}\|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}\|^{2}$ is essentially characterized as a linear combination of the top $m$ eigenvectors of $\mathbb{\Psi}^{\top}\mathbb{\Psi}$. Since a small gradient will prefer a small quantity of $\sum_{i=1}^{|\mathcal{V}|}\|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}\|^{2}$, the self-attention mechanism implicitly drives the weight matrices $\mathbb{W}_{Q}$ and $\mathbb{W}_{K}$ to align with the dominant eigen-directions of $\mathbb{\Psi}^{\top}\mathbb{\Psi}$. ∎
Appendix D Clustering in the Contextual Embedding Space
Clustering. We begin with the isotropy assesmment by performing clustering on the LLM representations in the contextual embedding space. There are various methods for performing cultering, such as $k$-means, DBSCAN Ester et al. [1996]. We select $K$-means clustering method because it is reasonably fast in high embedding dimensions (e.g., $d\geq 768$ for GPT2, ELMo, BERT etc.). We use the celebrated silhouette score analysis Rousseeuw [1987] to determine the number of clusters $|C|$ in the contextual embedding space. After performing $K$-means clustering, each observation $p$ (i.e., one of the $\mathbb{J}$ vector representations in $\mathcal{V}$) is assigned to one of $C$ clusters. For an observation $p$ assigned to the cluster $c\in C$, we compute the silhouette score as follows
$a(p)=\frac{1}{\left|C\right|-1}\sum\limits_{q\in C,p\neq q}\textrm{dist}(p,q);\qquad b(p)=\min_{\tilde{c}\neq c}\sum\limits_{q\in\tilde{c}}\textrm{dist}(p,q);\qquad s(p)=\frac{b(p)-a(p)}{\max b(p),a(p)},$
where $a(p)$ is the mean distance between an observation $p$ and the rest in the same cluster class $p$, while $b(p)$ measures the smallest mean distance from $p$-th observation to all observations in the other cluster class. After computing the silhouette scores $s(p)$ of all observations, a global score is computed by averaging the individual silhouette values, and the partition (with a specific number of clusters) of the largest average score is pronounced superior to other partitions with a different number of clusters. We select the best $|C|$ that belongs to the partition that scores highest among the other partitions.
Appendix E Dataset Description
Real Datsets.
Table 4: The complete list of datasets used for our quantitative and qualitative analysis. The table is divided into three sections, representing how the datasets were used for baseline models.
Dataset Domain Freq. Num. Series Series Length Prediction
min avg max Length ($H$)
Australian Electricity Energy 30min 5 230736 231052 232272 48
Car Parts Retail 1M 2674 51 51 51 12
Covid Deaths Healthcare 1D 266 212 212 212 30
Dominick Retail 1D 100014 201 296 399 8
Exchange Rate Finance 1B 8 7588 7588 7588 30
FRED-MD Economics 1M 107 728 728 728 12
Hospital Healthcare 1M 767 84 84 84 12
NN5 (Weekly) Finance 1W 111 113 113 113 8
Weather Nature 1D 3010 1332 14296 65981 30
Transportaion Signal Transport 1D 3010 1332 14296 65981 30
Synthetic (10 kernels) Numerical - 1000000 1024 1024 1024 64
Synthetic Datasets. We use KernelSynth Ansari et al. [2024], a method to generate synthetic dataset using Gaussian processes (GPs). KernelSynth allows generation of large, diverse datasets tailored to specific patterns or statistical properties, which is particularly useful when real-world data is scarce or incomplete. In this synthetic data generation process, the GPs are defined by a mean function, $\mu(t)$, and a positive definite kernel, $\kappa(x_{i},x_{j})$, which specifies a covariance function for variability across input pairs $(x_{i},x_{j})$. A kernel bank $\mathcal{K}$ (which consists of linear, RBF, and periodic kernels) is used to define diverse time series patterns. The final kernel $\tilde{\kappa}(x_{i},x_{j})$ is constructed by sampling and combining kernels from $\mathcal{K}$ using binary operations like $+$ and $\times$. Synthetic time series are generated by sampling from the GP prior, $GP(\mu(t)=0,\tilde{\kappa}(x_{i},x_{j}))$. The following algorithm presents the pseudocode for KernelSynth which essentially follows the approach in Ansari et al. [2024].
Algorithm 1 KernelSynth: Generating Synthetic Sequences via Gaussian Process Kernels
Input: Kernel bank $\mathcal{K}$, maximum kernels per time series $J=5$, and length of the time series $l_{\text{syn}}=1024$.
Output: A synthetic time series $\mathbf{x}_{1:l_{\text{syn}}}$.
1:$j\sim\mathcal{U}\{1,J\}$triangleright sample the number of kernels
[…]
3:$\kappa^{*}(t,t^{\prime})\leftarrow\kappa_{1}(t,t^{\prime})$
4:for $i\leftarrow 2$ to $j$ do
[…]
6:     $\kappa^{*}(t,t^{\prime})\leftarrow\kappa^{*}(t,t^{\prime})\star\kappa_{i}(t,t^{\prime})$triangleright compose kernels
7:end for
8:$\mathbf{x}_{1:l_{\text{syn}}}\sim\mathcal{GP}(0,\kappa^{*}(t,t^{\prime}))$triangleright draw a sample from the GP prior
9:return $\mathbf{x}_{1:l_{\text{syn}}}$
Appendix F Full Visualization of PCA plots for different models
The Variations in Chornos-T5’s hidden representations for different input context lengths is depicted in Figure 11
Figure 11: Variations in Chronos-T5’s hidden representations for different input context lengths for the same synthetic dataset “non-linear (Dataset 1)” : (a) Contextual embedding space for input context length $L=500$. (b) Contextual embedding space for input context length $L=100$.
F.1 Synthetic datasets

### v3

@@
 Then by projecting the representations into lower dimensions using these $m$ largest eigenvectors, we find the isotropy within the clusters in the contextual embedding space.
-(iii) Finally, we provide a comprehensive evaluation across $12$ real and $10$ synthetic
+(iii) Finally, we provide a comprehensive evaluation across $12$ real and $10$ synthetic time series datasets over $6$ different LLMs.
+Our experiments
 Figure 1: Time series tokenization.
-time series datasets over $6$ different LLMs.
+demonstrate that different characteristics of numerical data and model architectures could have different impacts on the isotropy.
 2 Problem Setup in Numerical Domains
@@
 Analysis settings.
-For illustrative purposes, in this section, we present two examples from a specific numerical domain, such as transport, and for a specific language model (i.e., GPT-2) that illustrate the conditions under which isotropy is preserved in LLM representations.
+For illustrative purposes, in this section, we present two examples from a specific numerical domain, such as transport, and for a specific language model (i.e., GPT-2) that illustrate the conditions under which isotropy is preserved
 Figure 2: Visualization of transport Dataset 1 (Left) and Dataset 2 (Right) labels for two extreme cases of velocities, i.e., $10$ km/hour and $100$ km/hour.
+in LLM representations.
 In Section 5, we provide comprehensive evaluations across $22$ different datasets from various numerical domains and other $5$ different language models with different architectures and tokenization techniques.
@@
 Datasets.
-We conduct a comprehensive evaluation using $12$ different real time series datasets from various numerical domains, including energy, nature, finance, healthcare, retail and transportation (data sources can be found in Table 4 of Appendix E).
+We conduct a comprehensive evaluation using $12$ different real time series datasets from various numerical domains, including energy, nature, finance, healthcare, retail and transportation.
+The sources of these open-source datasets along with their descriptions, including how each dataset is used across different LLM can be found in Table 4 of Appendix E.
 We also illustrate our findings using KernelSynth Ansari et al.
 (2024) (see Algorithm 1 in Appendix E for details), a method that generates $10$ more synthetic datasets via Gaussian processes in Section 5.
-In Section 4.1, we show how the LLM hidden representation varies for two different datasets from the same numerical domain and its impact on downstream task performance.
-So we select two different datasets from each numerical domain (for good and bad downstream task performance examples) for qualitative and quantitative analysis as shown in Table 3.
+As in Section 4.1, we select two different datasets from each numerical domain (as shown in Table 3) and then perform qualitative analysis with synthetic datasets and quantitative analysis with real datasets.
+The results of both of this analysis are provided in the next two sections.
+Figure 6: NMSE vs isotropy analysis for $10$ different synthetic datasets of $5$ different domains.
 5.1 Qualitative Analysis
 In this section, we analyze the time series forecasting by the baseline LLMs qualitatively.
-We focus on synthetically generated time series for a controlled analysis of different types of time series patterns which belong to $5$ different domains, such as linear, seasonality, trend, non-linear and Gaussian perturbation.
-Figure 6: NMSE vs isotropy analysis for $10$ different synthetic datasets of $5$ different domains.
-We are particularly interested in the isotropic measurement in the LLM’s last layer as it is related to the logits and probabilistic inference.
+We focus on synthetically generated time series for a controlled analysis of different types of time series patterns which belong to $5$ different domains, such as linear, seasonality, trend, non-linear and stochastic.
+We are particularly interested in the isotropic measurement in the LLM’s last layer as it is related to the logits and probabilistic inference as explained in Section 2.
 So all isotropic measure provided in this section is based on the last layer of the baselines.
 We begin by analyzing time series forecasting performance (i.e., NMSE) for different baselines and its relation with isotropy in Figure 6.
-For instance, in Figure 6 b, the inter-type cosine similarity in hidden embedding spaces of Chronos-T5 for seasonality (Dataset 1) is $-0.00007$ and for seasonality (Dataset 2) is $0.0047$.
-This indicates that stronger isotropy exists (i.e., inter-type cosine similarity value is close to 00) for Chronos-T5 for seasonality (Dataset 1) which preserves the structure in the Chronos-T5’s hidden representations and causes good downstream task performance.
-On the other hand, a weaker isotropy exists (i.e., inter-type cosine similarity value is far from 00) for Chronos-T5 in seasonality (Dataset 2), causing a lack of structure in Chronos-T5’s hidden representations, and thus leading to bad downstream task performance.
-Moreover, the NMSE for seasonality (Dataset 1) is lower for Chronos-T5, PatchTST and Lag-Llma, and higher for Chronos-Bolt and Morai, while the NMSE is higher for all the baselines except Morai for non-linear (Dataset 2).
-A similar analysis can also be observed for all other synthetic datasets and baselines in Figure 6.
-This shows that the same dataset from any particular domain may cause different forecasting performances for different baselines, as it generates different contextual embedding spaces for language models with different architectures and tokenization techniques (See Appendix F for full illustration).
+For instance, in Figure 6 b, we have (NMSE = $0.0000066$ and cosine similarity = $|-0.00076|$) for seasonality-Dataset 1 and (NMSE = $0.00012$ and cosine similarity = $0.0047$) for seasonality-Dataset 2 for Chronos-T5.
+This shows stronger isotropy exists (i.e., inter-type cosine similarity value is close to 00) in Chronos-T5’s embedding space for seasonality-Dataset 1 which preserves the structure in its hidden representations and causes good downstream task performance.
+On the other hand, a weaker isotropy exists (i.e., inter-type cosine similarity value is far from 00) in Chronos-T5’s embedding space for seasonality-Dataset 2, and thus causes a lack of structure in its hidden representations which leads to bad downstream task performance as compared to seasonality-Dataset 1.
+The NMSE and inter type cosine similarity may also vary for different language models and datasets.
+For example, in Figure 6 c, we can see that the the NMSE for trend-Dataset 1 is lower for PatchTST and Morai, while higher for Chronos-T5, Chronos-Bolt and Lag-Llma.
+In contrast, the NMSE for trend-Dataset 2 is lower for Chronos-T5, Chronos-Bolt, and Lag-Llma, while higher for PatchTST and Morai.
+A similar analysis can also be observed for other synthetic datasets and baselines in Figure 6 b, Figure 6 d, and Figure 6 e.
+This shows that any dataset from any particular domain may cause different forecasting performances for different baselines,
+Figure 7: NMSE vs isotropy comparison across different input context lengths for synthetic datasets.
+as it generates different hidden representations in contextual embedding spaces (see Appendix F for full visualization) for language models with different architectures and tokenization techniques.
 Next, we examine the influence of isotropy on forecasting performance in two important scenarios: a) different input context lengths, and b) different levels of noises in the input data.
-The first scenario is important as it provides an analysis on selecting proper input context length rather than selecting the length through random trials and errors.
+The first scenario is important as it provides an analysis that helps guide in selecting reasonable input context lengths rather than selecting the length through random trials and errors.
 The second scenario is important as it gives us ideas on how the level of noise in noisy data impacts performance, since the data in the real world is mostly noisy.
@@
 We first analyze the impacts of isotropy under varying input context lengths.
-We select Choronos-T5 as an example model and non-linear (Dataset 1) as an example dataset.
-In Figure 11 in Appendix F, we show how the hidden representations of Chronos-T5 vary for two different input context lengths, such as $L=500$ and $L=500$, for non-linear (Dataset 1), which gives an indication of having
-Figure 7: NMSE vs isotropy comparison across different input context lengths for synthetic datasets.
-different isotropic measures for different context lengths.
 In Figure 7, we compare the NMSE vs isotropy across two different input context lengths, $L=500$ and $L=100$, for different synthetic datasets.
+We use Choronos-T5 as an example model for this experiment.
 As can be seen from the figure, the isotropy values vary across different input context lengths and datasets.
-For instance, in seasonality (Dataset 1), we have (NMSE$=0.0000066$, cosine similarity$=-0.00076$) and (NMSE$=0.0793$, cosine similarity$=0.0011$) for $L=500$ and $L=100$, respectively.
+For instance, in seasonality-Dataset 1, we have (NMSE$=0.0000066$, cosine similarity$=|-0.00076|$) and (NMSE$=0.0793$, cosine similarity$=0.0011$) for $L=500$ and $L=100$, respectively.
 The decrease in isotropy significantly increases the NMSE for the input context length $L=100$.
-In contrast, in Linear (Dataset 2), the isotropy increases for the input context length $L=100$, which causes the decreases in NMSE for chornos-T5.
+In contrast, in linear-Dataset 2, we have (NMSE$=0.000025$, cosine similarity$=0.2474$) and (NMSE$=0.000009$, cosine similarity$=0.0644$) for $L=500$ and $L=100$, respectively.
+In this scenario, the isotropy increases for the input context length $L=100$, which causes the decrease in NMSE for chornos-T5.
+In Figure 11 in Appendix F, we show how the hidden representations of Chronos-T5 vary for two different input context lengths, such as $L=500$ and $L=500$, for non-linear-Dataset 1, which generates different isotropic measures for different input context lengths.
 In practice, the input context length is often selected randomly or through trial and error, which may cause higher forecasting errors for different datasets.
@@
 This insight helps guide improvements in forecasting performance by examining the isotropic properties of the contextual embedding space.
+Figure 8: NMSE vs isotropy comparison across different noise levels in synthetic datasets.
 Isotropy in varying noise levels in datasets.
@@
 From Figure 8, we can see consistently lower isotropy (i.e., inter-type cosine similarity far from 00) for all noisy synthetic datasets as compared to the datasets without noise.
-For instance, in trend (Dataset 2),
-Figure 8: NMSE vs isotropy comparison across different noise levels in synthetic datasets.
-we have (NMSE$=0.000024$, cosine similarity$=-0.00022$) and (NMSE$=0.0012$, cosine similarity$=0.0040$) for $\sigma=0$ and $\sigma=0.05$, respectively.
+For instance, in trend-Dataset 2, we have (NMSE$=0.000024$, cosine similarity$=|-0.00022|$) and (NMSE$=0.0012$, cosine similarity$=0.0040$) for $\sigma=0$ and $\sigma=0.05$, respectively.
 The decrease in isotropy significantly increases the NMSE for the noisy dataset.
-In practice, the environments of many real numerical domains, such as nature and energy, are noisy and dynamic.
-In such an environment, it is not always possible to measure the noise in real time and take the necessary steps to clean the input time series for better performance.
-However, it is always possible to measure the isotropy from LLM hidden representations, which can be used as a measure of noise in the input datasets, and thus, help to improve the forecasting performance.
+In practice, many real-world numerical domains—such as those in nature and energy—exhibit noisy and dynamic behavior.
+In these environments, it is often infeasible to measure noise in real time or to pre-process the input time series for improved performance.
+However, the isotropy in LLMs hidden representations can be readily measured, and thus, can be leveraged to enhance forecasting performance by identifying and mitigating the effects of noisy inputs in contextual embedding space.
 5.2 Quantitative Analysis
-In Figure 9, we analyze the time series forecasting performance for 12 real datasets with different baselines and its relation with isotropy.
-For instance, in Figure 9 e, the inter-type cosine similarity in hidden embedding spaces of Morai for retail (Dataset 1) is $0.002$ and for retail (Dataset 2) is $0.1931$.
-This indicates that stronger isotropy exists for Morai for retail (Dataset 1) which preserves the structure in the Morai’s hidden representations and causes good downstream task performance.
-On the other hand, a weaker isotropy exists for Morai in retail (Dataset 2),
+In this section, we present our main results on $12$ real datasets which belong to $6$ different numerical domains such as energy, nature, finance, healthcare, retail and transportation.
+As our qualitative analysis in Section 5.1, we select two different datasets from each numerical domain and the isotropy measure from LLM’s last layer to show the impact of isotropy on NMSE performance for different language models.
 Figure 9: NMSE vs isotropy analysis for $12$ different real datasets of $6$ different domains.
-causing a lack of structure in Morai’s hidden representations, and thus leading to bad downstream task performance.
-Moreover, the NMSE for retail (Dataset 2) is lower for all baselines except the Morai as compared to retail (Dataset 1), while the NMSE for energy (Dataset 1) is lower for all baselines except Lag-Llma as compared to energy (Dataset 2).
-A similar analysis can also be observed for all other real datasets and baselines in Figure 9.
-This again shows that a dataset may have different impacts on forecasting performances for different baselines, and isotropy can be used as a measure of how the dataset generates different contextual embedding spaces for different language models based on their architectures and tokenization techniques.
+In Figure 9, we analyze the time series forecasting performance of different baselines and its relation with isotropy for different real datasets.
+For instance, in Figure 9 e, we have (NMSE = $0.0061$ and cosine similarity = $0.0020$) for retail-Dataset 1 and (NMSE = $0.1255$ and cosine similarity = $0.1931$) for retail-Dataset 2 for Morai.
+This indicates the existence of stronger isotropy in Morai’s embedding space for retail-Dataset 1 which preserves the structure in its hidden representations and causes good downstream task performance.
+On the other hand, a weaker isotropy exists in Morai’s embedding space for retail-Dataset 2, and thus causes a lack of structure in its hidden representations which leads to bad downstream task performance as compared to retail-Dataset 1.
+The NMSE and inter type cosine similarity may also vary for different real datasets and language models.
+For example, in Figure 9 a, we can see that the NMSE for energy-Dataset 1 is lower for Chronos-T5, Chronos-Bolt, PatchTST and Morai, while higher Lag-Llma.
+In contrast, the NMSE for energy-Dataset 2 is lower for Morai while higher for other baselines.
+A similar analysis can also be observed for other synthetic datasets and baselines in Figure 9 b and Figure 9 c.
+This again shows that datasets from the same numerical domain can cause varying forecasting performance across different baselines, as they generate distinct hidden representations in contextual embedding spaces depending on the language model architecture and tokenization strategy.
+Figure 10: NMSE vs isotropy comparison across different input context lengths for real datasets.
 Finally, in Figure 10, we compare the NMSE vs isotropy for varying input context lengths to observe its impact on the real datasets.
 We select Lag-Llma as our example model.
-We compare the results for two different input context length: 1) the recommended input context length $L=144$ and the reduced input context length $L=96$.
-As can be seen from the figure, the violation in recommended input context length
-Figure 10: NMSE vs isotropy comparison across different input context lengths for real datasets.
-by reducing it from $L=144$ to $L=96$ not only decreases the NMSE performances, but also increases for some datasets.
-For instance, the inter-type cosine similarity values become close to 00, i.e., from $0.1091$ to $0.0012$ and from $0.2014$ to $0.0396$, respectively, for nature (Dataset 2) and finance (Dataset 1), which in turn improves the NMSE performances.
+We compare the results for two different input context lengths: 1) the recommended input context length $L=144$ and the reduced input context length $L=96$.
+As can be seen from the figure, the inter-type cosine similarity values become far from 00, i.e., from $0.0097$ to $0.0220$ for energy-Dataset 1 and from $0.0026$ to $0.0103$ for transport-Dataset 1, which in turn decreases the NMSE performances.
+On the other hand, the inter-type cosine similarity values become close to 00, i.e., from $0.1091$ to $0.0112$ for nature-Dataset 2 and from $0.2014$ to $0.0133$ for finance-Dataset 1, which in turn improves the NMSE performances.
+Thus, the variation in the recommended input context length may not only decrease the NMSE performances, but can also increase for some datasets.
 6 Conclusion and Limitations
@@
 $=|\mathbb{\Lambda}|_{2}\sum_{i=1}^{|\mathcal{V}|}\left(p_{i,i}+\frac{1}{2}\right)|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}p_{i,:}|^{2}+|\mathcal{V}|+|\mathbb{\Lambda}|_{2}\sum_{i\neq j}^{|\mathcal{V}|}p_{i,j}\left|\mathbb{\psi}_{j}-\sum_{q=1}^{|\mathcal{V}|}p_{i,q}\mathbb{\psi}_{q}\right|^{2}+\frac{|\mathbb{\Lambda}|_{2}}{2}\sum_{j=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}|^{2}$
-$=|\mathbb{\Lambda}|_{2}\sum_{i=1}^{|\mathcal{V}|}\left(p_{i,i}+\frac{1}{2}\right)|x_{i}-X^{\top}p_{i,:}|^{2}+|\mathcal{V}|+\Delta,$
-where $\Delta=|\mathbb{\Lambda}|_{2}\sum_{i\neq j}^{|\mathcal{V}|}p_{i,j}\left|\mathbb{\psi}_{j}-\sum_{q=1}^{|\mathcal{V}|}p_{i,q}\mathbb{\psi}_{q}\right|^{2}+\frac{|\mathbb{\Lambda}|_{2}}{2}\sum_{j=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}|^{2}.$
 The theorem below shows that $\mathbb{\Lambda}$ minimizing the objective $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}|^{2}$ contains the largest $m$ eigenvectors of the correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$ of input time series token embeddings where $m$ is the rank of $\mathbb{\Lambda}$.
@@
 The Variations in Chornos-T5’s hidden representations for different input context lengths is depicted in Figure 11
-Figure 11: Variations in Chronos-T5’s hidden representations for different input context lengths for the same synthetic dataset “non-linear (Dataset 1)” : (a) Contextual embedding space for input context length $L=500$.
+Figure 11: Variations in Chronos-T5’s hidden representations for different input context lengths for the same synthetic dataset non-linear-Dataset 1 : (a) Contextual embedding space for input context length $L=500$.
 (b) Contextual embedding space for input context length $L=100$.
-F.1 Synthetic datasets
+The full visualization of PCA plots of different models is provided below. We use the synthetic Dataset 1, and Dataset 2 from non-linear domain for illustration.

### v4

@@
 Abstract
-Recent studies have shown that vector representations of contextual embeddings learned by pre-trained large language models (LLMs) are effective in various downstream tasks in numerical domains.
+Vector representations of contextual embeddings learned by pre-trained large language models (LLMs) are effective in various downstream tasks in numerical domains such as time series forecasting.
 Despite their significant benefits, the tendency of LLMs to hallucinate in such domains can have severe consequences in applications such as energy, nature, finance, healthcare, retail and transportation, among others.
-To guarantee prediction reliability and accuracy in numerical domains, it is necessary to open the black-box and provide performance guarantees through explanation.
-However, there is little theoretical understanding of when pre-trained language models help solve numeric downstream tasks.
+To guarantee prediction reliability and accuracy in numerical domains, it is necessary to open the black box behind the LLM and provide performance guarantees through explanation.
+However, there is little theoretical understanding of when pre-trained language models help solve numerical downstream tasks.
 This paper seeks to bridge this gap by understanding when the next-word prediction capability of LLMs can be adapted to numerical domains through a novel analysis based on the concept of isotropy in the contextual embedding space.
-Specifically, we consider a log-linear model for LLMs in which numeric data can be predicted from its context through a network with softmax in the output layer of LLMs (i.e., language model head in self-attention).
-We demonstrate that, in order to achieve state-of-the-art performance in numerical domains, the hidden representations of the LLM embeddings must possess a structure that accounts for the shift-invariance of the softmax function.
-By formulating a gradient structure of self-attention in pre-trained models, we show how the isotropic property of LLM embeddings in contextual embedding space preserves the underlying structure of representations, thereby resolving the shift-invariance problem and providing a performance guarantee.
-Experiments show that different characteristics of numeric data and model architecture could have different impacts on isotropy.
+Specifically, a log-linear model for LLMs is considered in which numerical data can be predicted from its context through a network with softmax in the output layer of LLMs (i.e., language model head in self-attention).
+For this model, it is demonstrated that, in order to achieve state-of-the-art performance in numerical domains, the hidden representations of the LLM embeddings must possess a structure that accounts for the shift-invariance of the softmax function.
+By formulating a gradient structure of self-attention in pre-trained models, it is shown how the isotropic property of LLM embeddings in contextual embedding space preserves the underlying structure of representations, thereby resolving the shift-invariance problem and providing a performance guarantee.
+Experiments show that different characteristics of numerical data and model architectures have different impacts on isotropy, and this variability directly affects the performances.
 1 Introduction
-Large language models have demonstrated impressive success in adapting to various downstream tasks in numerical domains, such as finance Garza and Mergenthaler-Canseco (2023); Yu et al.
+Large language models (LLMs) have been proven to effective in adapting to various downstream tasks in numerical domains, such as finance Garza and Mergenthaler-Canseco (2023); Yu et al.
 (2023), energy Gao et al.
 (2024), climate science Jin et al.
-(2024), healthcare Wang and Zhang (2024), wireless communications Xu et al.
+(2024), healthcare Wang and Zhang (2024), transportation signals Xu et al.
 (2024), synthetic tabular generation Dinh et al.
@@
 (2024); Ansari et al.
-(2024) by adapting LLM to numerical domains.
-For many of these numeric downstream tasks, training a linear classifier on top of the hidden-layer representations generated by the pre-trained LLMs have already shown near state-of-the-art performance Jin et al.
+(2024) by adapting LLM to numerical domains that deal with time series forecasting.
+For many of these numerical downstream tasks, training a linear classifier on top of the hidden-layer representations generated by the pre-trained LLMs has been shown to achieve near state-of-the-art performance Jin et al.
 (2024); Ansari et al.
 (2024).
-However, these models in Gruver et al.
+However, the existing models in Gruver et al.
 (2024); Dooley et al.
@@
 (2024); Ansari et al.
-(2024) are treated as ‘black-box’ where numeric forecasts are controlled by complex nonlinear interactions between many parameters.
+(2024) are treated as a ‘black box’ where numerical forecasts are controlled by complex nonlinear interactions between many parameters.
 This makes it difficult to understand how models arrive at their predictions and makes it challenging for users to trust the model outputs.
-LLMs’ tendency to hallucinate can have serious consequences in critical numeric applications.
+When applied to critical numerical domain use cases, the tendency of LLMs to hallucinate can have serious and detrimental consequences.
 For example, prediction errors in fraud detection in finance can lead to huge financial losses and errors in protection onset of sepsis or cardiac arrest in healthcare can result in patient deaths.
-Thus, to guarantee prediction reliability and accuracy in numerical domains, it is necessary to open the black-box and provide performance guarantees through explaination.
+Thus, to guarantee prediction reliability and accuracy in numerical domains, it is necessary to understand the inner working of the so-called black box and provide performance guarantees through explanation.
 Although recent empirical studies Jin et al.
@@
 (2023); Liu et al.
-(2024) demonstrate the benefits of vector representations of embedding learned by LLMs in various numeric downstream tasks, there is little theoretical understanding of their empirical success.
+(2024) demonstrate the benefits of vector representations of embedding learned by LLMs in various numerical downstream tasks, there is little theoretical understanding of their empirical success.
 Thus, a fundamental question arises: “When (or how) can the next-word prediction capability of LLMs be effectively adapted to numerical domains?"
-The main contribution of this paper is to provide an approach to answer this question by exploiting the isotropic property of LLM hidden representations in the contextual embedding space.
+The main contribution of this paper is a novel approach for answering this question by exploiting the isotropic property of LLM hidden representations in the contextual embedding space.
 Isotropy refers to the geometric property wherein vector representations in the embedding space are uniformly distributed in all directions, a characteristic critical for maintaining the expressiveness of the embedding space Arora et al.
@@
 To achieve state-of-the-art performance in numerical domains, we show that the hidden representations of LLMs must exhibit a structured form in contextual embedding space that accounts for the shift-invariance of the softmax function (i.e., the softmax output remains unchanged when all logits are shifted by a constant).
-Without such structure, the model can shift the logits while keeping the training loss unchanged, thereby leaving the logits ineffective for numeric downstream tasks.
-By formulating a gradient structure of self-attention in pre-trained models, we show how the isotropic property of LLM embeddings in contextual embedding space preserves the underlying structure of representations, thereby resolving the shift-invariance problem of the softmax function.
-Our key contributions include: (i) We consider a log-linear model for LLMs and demonstrate theoretically why hidden representations must exhibit structure to address the shift-invariance problem of the softmax function.
-(ii) We take a deeper look into the hidden representations of pre-trained models and show how isotropy preserves the structural integrity of representations.
-In particular, we derive an upper bound for the Jacobian matrix which collects all first-order partial derivatives of self-attention with respect to the input pattern and show that $m$ largest eigenvectors of the LLM hidden representations minimize the gradient norm of self-attention.
-Then by projecting the representations into lower dimensions using these $m$ largest eigenvectors, we find the isotropy within the clusters in the contextual embedding space.
-(iii) Finally, we provide a comprehensive evaluation across $12$ real and $10$ synthetic time series datasets over $6$ different LLMs.
-Our experiments
-Figure 1: Time series tokenization.
-demonstrate that different characteristics of numerical data and model architectures could have different impacts on the isotropy.
+Without such structure, the model can shift the logits while keeping the training loss unchanged, thereby leaving the logits ineffective for numerical downstream tasks.
+By formulating a gradient structure of self-attention in pre-trained models, we show how the isotropic property of LLM embeddings in the contextual embedding space preserves the underlying structure of representations, thereby resolving the shift-invariance problem of the softmax function.
+In a nutshell, our key contributions include:
+  • We consider a log-linear model for LLMs and demonstrate theoretically why hidden representations must exhibit structure to address the shift-invariance problem of the softmax function.
+  • We take a deeper look into the hidden representations of pre-trained models and show how isotropy preserves the structural integrity of representations.
+In particular, we derive an upper bound for the Jacobian matrix which collects all first-order partial derivatives of self-attention with respect to the input pattern and show that the $m$ largest eigenvectors of the LLM hidden representations minimize the gradient norm of self-attention.
+Then, by projecting the representations into lower dimensions using these $m$ largest eigenvectors, we find the isotropy within the clusters in the contextual embedding space.
+  • Finally, we provide a comprehensive evaluation across $12$ real and $10$ synthetic time series datasets over $5$ different LLMs.
+Our experiments demonstrate that the isotropy of LLM hidden representations varies significantly based on the input data characteristics (i.e., domain, context length and noise level) and model design choices (i.e., tokenization techniques and architecture), which in turn strongly influences forecasting performance in numerical domains.
 2 Problem Setup in Numerical Domains
 Time Series Tokens and Similarity Measure.
-Similar to next-word prediction by LLMs, the next-value prediction in the numerical domain can be modeled by time series forecasting techniques Jin et al.
+Similar to next-word prediction by LLMs, the next-value prediction in the numerical domain can be modeled by time series forecasting techniques which are widely adopted in the machine learning literature Jin et al.
 (2024); Ansari et al.
-(2024) which are widely adopted in the machine learning literature.
-Formally, given a time series $\textbf{x}_{1:T+L}=[x_{1},\ldots,x_{T},\ldots,x_{T+L}]$, where the first $T$ time instances give the historical context, the next $L$ time instances constitute the forecast region, and $x_{t}\in\mathbb{R}$ is the observation of each time instance, we are interested in predicting the joint distribution of next $L$ time instances, $p(\textbf{x}_{T+1:T+L}|\textbf{x}_{1:T})$.
+(2024).
+Formally, given a time series $\textbf{x}_{1:T+L}=[x_{1},\ldots,x_{T},\ldots,x_{T+L}]$,
+Figure 1: Time series tokenization.
+where the first $T$ time instances give the historical context, the next $L$ time instances constitute the forecast region, and $x_{t}\in\mathbb{R}$ is the observation of each time instance, we are interested in predicting the joint distribution of the next $L$ time instances, $p(\textbf{x}_{T+1:T+L}|\textbf{x}_{1:T})$.
 Since, the pre-trained models operate on tokens from a finite vocabulary, using them for time series data requires mapping the observations to a finite set of tokens.
-Based on different numeric applications and LLM architectures, various tokenization techniques, e.g., quantization and scaling Ansari et al.
+Based on different numerical applications and LLM architectures, various tokenization techniques, e.g., quantization and scaling Ansari et al.
 (2024); Rasul et al.
@@
 (2024); Nie et al.
-(2023), and adaptation of language model tokenizer in numeric domians Gruver et al.
+(2023), and adaptation of language model tokenizer in numerical domians Gruver et al.
 (2024); Dooley et al.
@@
 Here, different contexts in the time series sequences yield different LLM embeddings of $k_{i}$.
-By constructing $\sum_{k}|\tilde{\mathbb{\Psi}}(k)|=|\mathcal{V}|$, we define the inter-token cosine similarity as,
+By constructing $\sum_{k}|\tilde{\mathbb{\Psi}}(k)|=|\mathcal{V}|$, we define the inter-token cosine similarity as:
 $\zeta_{\textrm{cos}}\triangleq\mathbb{E}_{i\neq j}[\cos(\mathbb{\psi}(k_{i}),\mathbb{\psi}(k_{j}))],\vspace{-0.9in}$ (1)
@@
 The inter-token cosine similarity metric describes the similarity between different tokens based on the contexts.
-For ease of reading, we express $T+l$ as $T_{l}$ and $T+l+1$ as $T_{l+1}$ for the rest of the paper.
+For notational simplicity, we express $T+l$ as $T_{l}$ and $T+l+1$ as $T_{l+1}$ hereinafter.
 Model.
 We consider a general pre-trained model for numerical data and open the black box of the pre-trained model.
-Specifically, we assume that the observation probability of $k_{T_{l+1}}$ given $\mathbb{k}_{1:T_{l}}$ satisfies the log-linear model Arora et al.
+Specifically, we assume that the observation probability of $k_{T_{l+1}}$ given $\mathbb{k}_{1:T_{l}}$ satisfies the log-linear model as in Arora et al.
 (2016)
@@
 (2021).
-To define the numeric downstream task, let $z^{*}_{i}(k,l)\!\!:=\!\!\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle$ be the $i$-th logit of the ground-truth model, and assume that the numeric downstream tasks are defined by a function of the logits, i.e., $f^{*}(\mathbb{z^{*}})$.
+To define the numerical downstream task, let $z^{*}_{i}(k,l)\!\!:=\!\!\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle$ be the $i$-th logit of the ground-truth model, and assume that the numerical downstream tasks are defined by a function of the logits, i.e., $f^{*}(\mathbb{z^{*}})$.
 Also let $Z^{*}(k,l)\!\!=\!\!\sum_{i=1}^{N}\exp(z^{*}_{i}(k,l))\!=\!\sum_{i=1}^{|\mathcal{V}|}\exp(\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle)$ be the partition function Arora et al.
@@
 Then, the normalized ground-truth model $\forall i\!\in\!\mathcal{V}$ is given by
-$p(k_{T_{l+1}}\!\!=i\,\,|\,\,\mathbb{k}_{1:T_{l}})$ $=\!\frac{\exp(\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle)}{\sum_{i=1}^{|\mathcal{V}|}\exp(\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle)}=\frac{\exp(z^{*}_{i}(k,l))}{Z^{*}(k,l)}$
-Since we do not know the ground-truth model in reality, we do not have access to the ground-truth model $\mathbb{\psi}^{*}(k_{i})$ and $\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}})$.
+$p(k_{T_{l+1}}\!\!=i\,\,|\,\,\mathbb{k}_{1:T_{l}})$ $=\!\frac{\exp(\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle)}{\sum_{i=1}^{|\mathcal{V}|}\exp(\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle)}=\frac{\exp(z^{*}_{i}(k,l))}{Z^{*}(k,l)}.$
+Since we do not know the ground-truth model in reality, we do not have access to the ground-truth model components $\mathbb{\psi}^{*}(k_{i})$ and $\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}})$.
 Instead, we only have access to the student model $\mathbb{\psi}(k_{i})$ and $\mathbb{\mathbb{\psi}}(\mathbb{k}_{1:T_{l}})$ that aims to achieve low pre-training loss.
@@
 (2023).
-Then, it is more reasonable to model the numeric downstream task as
+Then, it is more reasonable to model the numerical downstream task as
 $f^{*}(k,l)$ $=\sum_{i=1}^{|\mathcal{V}|}a^{*}_{i}\sigma(z^{*}_{i}(k,l)-b^{*}_{i})=\sum_{i=1}^{|\mathcal{V}|}a^{*}_{i}\sigma(\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle-b^{*}_{i}),\vspace{-1.5in}$
 where $\sigma$ is the ReLU function and $b^{*}_{j}$ denotes the threshold for the logits.
-The numeric downstream task only considers the logits that are above the threshold, and thus ignores all the entries with very small probabilities.
+The numerical downstream task only considers the logits that are above the threshold, and thus ignores all the entries with very small probabilities.
+Despite the empirical success of language models in numerical domains, there remains a fundamental gap in understanding when and why these models generalize reliably to numeric downstream tasks such as time series forecasting across different numerical domains.
+A key challenge lies in the mismatch between training and inference behavior, i.e., good training performance does not always guarantee robust performance at inference time.
+To address this challenge, we propose a novel theoretical framework grounded in the isotropic property of the contextual embedding space.
+We show that the presence of strong isotropy in LLM hidden representations stabilizes the partition function, effectively resolving the softmax shift-invariance problem and leading to reliable inference performance.
+The next section formalizes this insight and provides theoretical justification for using isotropy as a key indicator of model reliability in numerical settings.
 3 The Role of Isotropy in Adapting LLMs to Numerical Data
-As previously discussed in Section 2, we consider LLM networks whose last layer is usually a softmax layer and the numeric downstream task is determined by the function of the logits.
-The underlying relation between the logits and softmax function determines the performance of the numeric downstream tasks.
+As discussed in Section 2, we consider LLM networks whose last layer is usually a softmax layer and the numerical downstream task is determined by the function of the logits.
+The underlying relation between the logits and softmax function determines the performance of the numerical downstream tasks.
 However, the softmax function is shift-invariant, that is, the output of the softmax function remains unchanged when all logits are shifted by a constant.
-Since we do not have any control over the logit shift of the student model on unseen data, good performance during training does not necessarily provide any performance guarantee for the numeric downstream task on unseen scenarios.
-This can be formalized in the following theorem and the proof is provided in Appendix A.
-Theorem 3.1.
+Since we do not have any control over the logit shift of the student model on unseen data, good performance during training does not necessarily provide any performance guarantee for the numerical downstream task on unseen scenarios.
+The lack of perfromance guarantee under uncontrolled logit shifts on unseed data can be formalized in the following theorem:
+Theorem 1.
 Let the logits of the ground-truth model be bounded.
 Then for any $f^{*}(k,l)$, there exists a set of functions $\{\hat{z}_{i}(k,l)\}_{i=1}^{|\mathcal{V}|}$ such that for all $k$ and $T_{l+1}$, the predictive distribution of the student model $\hat{p}\,(k_{T_{l+1}}\,|\,\mathbb{k}_{1:T_{l}})$ matches that of ground-truth model $p^{*}(k_{T_{l+1}}\,|\,\mathbb{k}_{1:T_{l}})$ and $\hat{f}(k,l)=0$.
-In other words, there exists a student model with the same pre-training loss as the ground-truth model, but its logits are ineffective for the numeric downstream tasks.
-Theorem 3.1 demonstrates that without any structure in the hidden representations of LLM embeddings, the student model is able to shift the logits for any sample while keeping the pre-training loss unchanged and leaving logits ineffective for the numeric downstream tasks.
-Consequently, a theoretical guarantee for the numeric downstream task performance needs structure in the LLM representations learned by the pre-trained model.
-One way to prevent the shift-invariance problem from influencing the performance of the numeric downstream tasks is to keep the partition function stable.
+In other words, there exists a student model with the same pre-training loss as the ground-truth model, but its logits are ineffective for the numerical downstream tasks.
+Proof.
+The proof is provided in Appendix A. ∎
+Theorem 1 shows that without any structure in the hidden representations of LLM embeddings, student model can shift the logits for any sample while keeping the pre-training loss unchanged and leaving logits ineffective for the numerical downstream tasks.
+Consequently, a theoretical guarantee for numerical downstream task performance will require structure in the LLM representations learned by the pre-trained model.
+In this paper, we make an observation that to prevent the shift-invariance problem from influencing the performance of the numerical downstream tasks, it is necessary to keep the partition function stable.
 Let $\mathbb{\Psi}\!=\!(\mathbb{\psi}_{1}(k),\ldots,\mathbb{\psi}_{|\mathcal{V}|}(k))^{\top}\!\in\!\mathbb{R}^{|\mathcal{V}|\times D}$ be the hidden representations of input time series sequence.
-Then the stability of the partition function can be assessed through the isotropy in the contextual embedding space Arora et al.
+Then, the stability of the partition function can be assessed through the isotropy in the contextual embedding space Arora et al.
 (2016); Mu and Viswanath (2018) as follows
@@
 If the partition function remains stable for different samples, the logits can be solely determined by the probabilities, thereby resolving the shift-invariance problem of the softmax function.
+Building on this theoretical foundation, we now turn to the following empirical question: “How can we measure and interpret isotropy in practice, and how does it relate to generalization across numerical domains?".
+Motivated by Theorem 3.1 and the need for structural constraints in LLM representations, we analyze the effective dimensionality and cluster organization of LLM’s hidden representations in the contextual embedding space.
+These analyses reveal how isotropy manifests in pre-trained LLMs and how its presence correlates with the model’s ability to generalize to time series forecasting across different numerical domains, and hence, provides a performance guarantee.
+Section 4 introduces methods for quantifying this internal structural integrity using spectral alignment mechanism, principal component analysis (PCA) and cluster-based isotropy metric, and consequently, linking theoretical reliability (i.e., performance guarantee) to empirical generalizability.
 4 Study of isotropy in LLM hidden representations
 Analysis Settings.
-For illustrative purposes, in this section, we present two examples from a specific numerical domain, such as transport, and for a specific language model (i.e., GPT-2) that illustrate the conditions under which isotropy is preserved
-Figure 2: Visualization of transport Dataset 1 (Left) and Dataset 2 (Right) labels for two extreme cases of velocities, i.e., $10$ km/hour and $100$ km/hour.
-in LLM representations.
-In Section 5, we provide comprehensive evaluations across $22$ different datasets from various numerical domains and other $5$ different language models with different architectures and tokenization techniques.
-For this specific analysis, we select two datasets from transport as they are dynamic, noisy, time-varying, and thus hold all primary characteristics of the numerical data across various domains.
-Specifically, we use two ways of sharing signals which are typically used for signal transmission in transportation settings, as shown in Figure 2.
-We call these two datasets as “Dataset $1$” and “Dataset $2$”.
-The downstream task here is to predict the transport signalling property over wireless channel using LLM, where Dataset $1$ causes good downstream performance (i.e., near optimal signal prediction), while the Dataset $2$ causes bad downstream task performance (i.e., high error in signal prediction).
-We use NMSE as a performance metric for the numeric downstream task because it is widely used for signal prediction.
-We deploy the first six layers of GPT2 Radford et al.
-(2019) and use the datasets and simulation setups from Liu et al.
-(2024), which are the standard settings for time series forecasting.
-We predict $L=4$ future signaling properties based on the historical $T=16$ signal properties through time series forecasting using GPT2.
-The training and validation dataset contains $8,000$ and $1,000$ samples, respectively, with user velocities uniformly distributed between $10$ km/hour and $100$ km/hour.
-The test dataset contains ten velocities ranging from $10$ km/hour to $100$ km/hour, with $1,024$ samples for each velocity.
+For this study, we consider five different language models including Chronos-T5 Ansari et al.
+(2024), Chronos-Bolt [^1] https://huggingface.co/autogluon/chronos-bolt-base, PatchTST Nie et al.
+(2023), Moirai-1.0-R Woo et al.
+(2024), and Lag-Llama Rasul et al.
+(2024).
+For illustration, we randomly select a real dataset (i.e., finance-Dataset 1) from a broader collection of 22 numerical datasets that we use in this paper since we see similar results with all of these datasets.
+The details of these models and datasets could be found in Section 5.
 Table 1: The effective dimension $d(0.8)$
-Layer 1 2 3 4 5 6
-Dataset 1 4 4 4 4 4 4
-Dataset 2 1 1 1 1 1 1
-Effective Dimensions.
+Layer 1 2 3 4 5 6 7 8 9 10 11 12
+Chronos-T5 4 4 4 4 4 4 4 4 4 4 4 4
+Chronos-Bolt 1 1 1 1 1 1 1 1 1 1 1 1
+PatchTST 1 1
+Moirai 1 1 1 1 1 1
+Lag-Llma 2 2 2 2 2 2 2 2
+4.1 Effective Dimensions
 In each layer of each model, we start with a data matrix $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times D}$, where $|\mathcal{V}|$ represents the number of tokens in the input time series sequence, and $D$ corresponds to the embedding dimension.
 We apply PCA to reduce the dimensionality from $D$ to $m$ i.e., $\mathbf{\tilde{A}}\in\mathbb{R}^{|\mathcal{V}|\times m}$.
-Then the fraction of variance captured by the reduced representation is given by: $r_{m}=\frac{\sum_{i=0}^{m-1}\sigma_{i}}{\sum_{i=0}^{D-1}\sigma_{i}}$ where $\sigma_{i}$ denotes the $i$-th largest eigenvalue of the covariance matrix of $\mathbf{A}$.
+Then, the fraction of variance captured by the reduced representation is given by: $r_{m}=\frac{\sum_{i=0}^{m-1}\sigma_{i}}{\sum_{i=0}^{D-1}\sigma_{i}}$ where $\sigma_{i}$ denotes the $i$-th largest eigenvalue of the covariance matrix of $\mathbf{A}$.
 We define the $\epsilon$-effective dimension as $d(\epsilon)\triangleq\arg\min_{m}r_{m}\geq\epsilon$.
-For instance, if $d(0.8)=3$, it means that three principal dimensions retain $80$% of the variance.
+For instance, if $d(0.8)=3$, then three principal dimensions retain $80$% of the variance.
 A higher $d$ suggests a more isotropic space Cai et al.
@@
 Table 1 presents the values of $d(0.8)$ for different layers and models.
-Surprisingly, GPT-2 has so few effective dimensions, with $d(0.8)=4$ for Dataset $1$ and $d(0.8)=4$ for Dataset $1$, for layers 1 through 6, as compared to its original embedding dimensions $D=768$.
-The reduced dimensionality suggests that GPT-2’s embeddings lie in a subspace defined by a very narrow cone Ethayarajh (2019), and hence, there would be hardly any model capacity.
-Surprisingly, these language models are as successful as they are in numerical domains, given that most of their embedding vectors are as similar to one another as they are.
-These observations motivate us to look deeper into the contextual embedding space.
-4.1 Clusters in the Contextual Embedding Space
+Surprisingly, all of these models have very small effective dimensions as compared to original embedding dimensions.
+For instance, Chronos-Bolt has very small effective dimensions, with $d(0.8)=1$ for layers 1 through 12, as compared to its original embedding dimensions $D=512$.
+The small effective dimensionality is another way of telling that that Chronos-Bolts’s embedding vectors lie in a subspace defined by a very narrow cone Ethayarajh (2019), and consequently, their inter-token cosine similarity is large.
+If all the embedding vectors lie on a $1$-dimensional line, the inter-token cosine similarity would be close to $1$, and there would be hardly any model capacity.
+Surprisingly, despite having such low effective dimensionality, these language models still perform well in numerical domains.
+This counterintuitive result motivate us to look deeper into the contextual embedding space.
+4.2 Spectral Alignment for Generalization in Numerical Settings
 Let $G(\mathbb{\Psi})\!=\!(g_{1}(\mathbb{\Psi}),\ldots,g_{|\mathcal{V}|}(\mathbb{\Psi}))^{\top}\!:\mathbb{R}^{|\mathcal{V}|\times D}\mapsto\mathbb{R}^{|\mathcal{V}|\times D}$ be the function for self-attention, i.e., $g_{i}(\mathbb{\Psi})=\text{softmax}(\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\Psi}^{\top})\mathbb{\Psi}$, where $\mathbb{\Lambda}\!=\!\mathbb{W}_{Q}\mathbb{W}_{K}^{\top}\in\mathbb{R}^{D\times D}$, and $\mathbb{W}_{Q}\!\!\in\!\!\mathbb{R}^{D\times m}$, $\mathbb{W}_{K}\!\!\in\!\!\mathbb{R}^{D\times m}$ are the parameter matrices for the query and key matrices of self-attention.
-The lemma below contributes to understanding why the isotropic property of pre-trained LLMs help to generalize to numerical domains.
-The proof of this lemma closely aligns with the approach in Kim et al.
-(2021).
+The lemma below provides insights into how the isotropic property of pre-trained LLMs enables generalization in numerical domains.
 The proof of this lemma follows the analysis in Kim et al.
 (2021) is provided in Appendix B for completeness.
-Lemma 4.1.
+Lemma 1.
 Consider the Jacobian matrix $\mathbb{J}=\left[\frac{\partial g_{i}(\mathbb{\Psi})}{\partial\mathbb{\psi}_{j}}\right]_{i,j=1}^{|\mathcal{V}|}$, which represents the gradient of the self-attention mapping $G(\mathbb{\Psi})$ with respect to the input time series token embeddings.
 Then the spectral norm of $\mathbb{J}$ satisfies $\|\mathbb{J}\|_{2}\leq|\mathbb{\Lambda}|_{2}\sum_{i=1}^{|\mathcal{V}|}\left(p_{i,i}+\frac{1}{2}\right)\left|\mathbb{\psi}_{i}-\sum_{j=1}^{|\mathcal{V}|}p_{i,j}\mathbb{\psi}_{j}\right|^{2}+\Delta$, where the residual term $\Delta$ is given by $\Delta=|\mathbb{\Lambda}|_{2}\sum_{i\neq j}^{|\mathcal{V}|}p_{i,j}\left|\mathbb{\psi}_{j}-\sum_{q=1}^{|\mathcal{V}|}p_{i,q}\mathbb{\psi}_{q}\right|^{2}+\frac{|\mathbb{\Lambda}|_{2}}{2}\sum_{j=1}^{|\mathcal{V}|}|\mathbb{\psi}_{j}|^{2}$, and the attention weights $p_{i,j}$ are defined as $p_{i,j}=\frac{\exp(\mathbb{\psi}_{i}^{\top}\mathbb{\Lambda}\mathbb{\psi}_{j})}{\sum_{k=1}^{|\mathcal{V}|}\exp(\mathbb{\psi}_{i}^{\top}\mathbb{\Lambda}\mathbb{\psi}_{k})}$.
-From Lemma 4.1, we can see that to minimize the norm of the gradient $\|\mathbb{J}\|_{2}$, we essentially need to make $\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\sum_{j=1}^{|\mathcal{V}|}p_{i,j}\mathbb{\psi}_{j}\right|^{2}$ small.
+From Lemma 1, we can see that, in onder to minimize the norm of the gradient $\|\mathbb{J}\|_{2}$, we essentially need to make $\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\sum_{j=1}^{|\mathcal{V}|}p_{i,j}\mathbb{\psi}_{j}\right|^{2}$ small.
 When $\mathbb{\Lambda}$ is small and all the input time series token embeddings are centered at the origin, $\sum_{i=1}^{|\mathcal{V}|}\mathbb{\psi}_{i}\!=\!0$, we have $\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}p_{i,:}\right|^{2}\approx\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}\right|^{2}$ (see Appendix B).
-The theorem below shows that $\mathbb{\Lambda}$ minimizes the objective $\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}\right|^{2}$ and contains the $m$ largest eigenvectors of correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$ of time series token embeddings, where $m$ is the rank of $\mathbb{\Lambda}$.
-The proof of Theorem 4.1 is provided in Appendix C.
-Theorem 4.1.
+Next, we prove that $\mathbb{\Lambda}$ minimizes the objective $\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}\right|^{2}$ and contains the $m$ largest eigenvectors of correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$ of time series token embeddings, where $m$ is the rank of $\mathbb{\Lambda}$.
+Theorem 2.
 Let the eigenvalues of the correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$ be ordered as $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{D}$, and let $\gamma_{i}\in\mathbb{R}^{D}$ for $i=1,\ldots,D$ denote their associated eigenvectors.
 Then, the matrix $\mathbb{\Lambda}^{*}$ that minimizes the quantity $\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}\right|^{2}$ has the optimal form $\mathbb{\Lambda}=\sum_{i=1}^{m}\frac{1}{\lambda_{i}}\gamma_{i}\gamma_{i}^{\top}.$
-Theorem 4.1 shows that the self-attention learns to perform a function for the numeric downstream tasks through training, which are closely related to the $m$ largest eigenvectors of LLM hidden representations.
-In other words, the self-attention mechanism effectively projects input time series tokens onto a low-dimensional contextual embedding space defined by the top eigenvectors of correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$.
-Figure 3: Variations in GPT-2’s hidden representation for different datasets from the same domain: (a) PCA plot of contextual embedding space for transport Dataset 1.
-(b) PCA plot of contextual embedding space for transport Dataset 2.
-Motivated by the findings from Lemma 4.1 and Theorem 4.1, we project the models’ hidden representations of into a lower-dimensional space by utilizing the $m\!=\!3$ largest eigenvectors through PCA, as shown in Figure 3.
+Proof.
+The proof of Theorem 2 is provided in Appendix C. ∎
+Theorem 2 shows that the self-attention mechanism effectively projects input time series tokens onto a low-dimensional contextual embedding space defined by the top eigenvectors of the correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$.
+This result reveals that the self-attention mechanism in LLMs implicitly aligns with the dominant directions (i.e., top eigenvectors) of the contextual embedding space, and hence, suggesting that isotropy is not just a geometric artifact but a learned structural property that supports effective generalization to numerical downstream tasks.
+While the self-attention aligns input representations with the dominant eigenvectors of the embedding space, the alignment may vary across different subregions of the contextual embedding space due to variations in the input sequences, token types, or contextual patterns.
+As a result, the degree of isotropy may differ across subregions of the contextual embedding space, which motivates the need to assess isotropy at a local (i.e., cluster) level rather than relying solely on a global metric.
+The next section explores these local structural patterns and examines the geometry of the hidden representations through principal component analysis (PCA), which helps reveal how variance is distributed across embedding dimensions.
+Figure 2: Isolated or slightly overlapping cluster islands exist in the contextual embedding space for all models. For brevity, we only show a few representative middle layers from each model.
+4.3 Clusters in the Contextual Embedding Space
+Motivated by the results of Lemma 1 and Theorem 2, this section investigates local structural patterns by projecting the LLMs’ hidden representations into a lower-dimensional space using the top $m\!=\!3$ eigenvectors via PCA, as shown in Figure 2.
 The three axes of the figure represent the first three principal components of the covariance matrix of LLM representations of each layer.
-For instance, in layer $6$, the first three principal components account for $76$% and $92$% of the total variance for Dataset $1$ and Dataset $2$, respectively.
-From Figure 3 a and Figure 3 b, we can see that there are disconnected or slightly overlapped islands that are far away from each other through layer 3 to layer 6.
-Note that the first principal dimension value spans from 00 to $2000$, significantly wider than the other $2$ dimensions, and dominates the total variance.
-A similar analogy can also be observed for transport Dataset 2 in Figure 3 b.
+For instance, in Figure 2b and 2d, the first three principal components account for $94$% of the total variance in layer 8 of Lag-Llama and $83$% in layer 2 of Moirai.
+From Figure 2 a, 2 b, 2 c and 2 d, we can see that there are disconnected or slightly overlapping islands that are far away from each other.
 In (1), the space isotropy is measured on pairs of arbitrary time series token representations, which could reside in two disconnected clusters.
 However, given that the variance is dominated by distances between clusters, such estimation would be biased by the inter-cluster distances.
-Hence, it is more meaningful to consider a per-cluster (i.e., local) investigation rather than a global estimate.
+Hence, it is more reasonable to consider a per-cluster (i.e., local) investigation rather than a global estimate.
+Isotropy within Clusters.
 We start by performing clustering on the LLM representations in the contextual embedding space.
-There are various methods for performing clustering, such as $k$-means, DBSCAN Ester et al.
+There are various methods for performing clustering, such as $K$-means and DBSCAN algorithm Ester et al.
 (1996).
@@
 A value close to 0 indicates strong isotropy and ensures the existence of structure in the LLM representations.
-Isotropy in Dataset 1.
-In this section, we provide an example of Dataset $1$ where a high prediction accuracy (i.e., low NMSE) is achieved by GPT2 based model.
-For instance, in Figure 4a, we compare the NMSE performance of our GPT2 based LLM with non-language models for different user velocities.
-From Figure 4 a, we can observe that the NMSE performance of all models gradually increased along with the increase in user velocity.
-This is because, with the increase in velocity, the transport signal characteristics rapidly change within a very short coherence time, resulting in increased prediction difficulty for the prediction model.
-The GPT2 based model consistently outperforms other baselines and demonstrates its high prediction accuracy.
-Figure 4: (a) LLM outperforms all other baselines for all of the ten different velocities for Dataset 1.
-(b) Inter-type cosine similarities for Dataset $1$ with different velocities.
-$\zeta\textquoteright_{\textrm{cos}}$ are close to zero for all the layers, including layer 6, indicating that nearly perfect isotropy exists in the LLM embedding space for the Dataset $1$, which preserves the structure in the LLM’s hidden representations and causes the high prediction accuracy.
-For illustrative purposes, we pick three user velocities: $10$ km/hour, $50$ km/hour, and $100$ km/hour, for isotropy assessment of Dataset $1$.
-The GPT2 based model achieves good NMSE performance for all of these three velocities, as shown in Figure 4a.
-We apply inter-type cosine similarity $\zeta^{\prime}_{\textrm{cos}}$ in (6) to measure the isotropy in GPT2 embedding space.
-From Figure 4b, we can see that the GPT2 has consistent near-zero cosine similarity values for all layers, including layer 6.
-This indicates that nearly perfect isotropy exists in the GPT2 embedding space for the Dataset $1$, which preserves the structure in the GPT2’s hidden representations and causes good downstream task performance.
-Figure 5: (a) The NMSE performance of the LLM based model for Dataset 2 deteriorates significantly compared to Dataset 1.
-(b) Inter-type cosine similarities for Dataset $2$ for different velocities.
-Higher $\zeta_{\textrm{cos}}$ values indicate a weak isotropy (i.e., anisotropy) exists in the LLM embedding space which causes a lack of structure in the hidden representations, yielding poor prediction accuracy.
-Isotropy in Dataset 2.
-In this section, we provide an example of Dataset $2$ where a poor prediction accuracy (i.e., high NMSE) is achieved by GPT2 based model.
-As shown in Figure 5 a, the NMSE performance fluctuates randomly for different velocities, while the NMSE performance for Dataset $1$ is gradually increasing with increase in the velocities.
-The NMSE performance for Dataset $2$ deteriorates significantly compared to the Dataset $1$.
-As before, with the three user velocities, the NMSE performance for Dataset $2$ for all of these velocities is worse as compared to Dataset $1$, as shown in Figure 5 a.
-From Figure 5 b, we can observe a weak isotropy (i.e., anisotropy) in the LLM embedding space for Dataset $2$, causing a lack of structure in the GPT2 hidden representations, and thus leading to bad downstream performance.
+To put it in a nutshell, this section provides a theoretical foundation showing that self-attention projects input tokens onto a low-dimensional subspace aligned with the dominant eigenvectors of the embedding correlation matrix.
+This alignment induces isotropy in LLM hidden representations, stabilizing the partition function and preserving the structure needed for reliable numerical downstream task performances.
+In Section 5, we extend this analysis by empirically evaluating how isotropy in different language models’ hidden representations correlates with time series forecasting performances across a wide range of numerical datasets, varying context lengths, and noise levels.
 5 Experiments
 Baselines.
-We consider popular pre-trained LLMs as the baselines for numeric downstream tasks, including PatchTST Nie et al.
-(2023), Lag-Llama Rasul et al.
+We consider popular pre-trained LLMs as the baselines for numerical downstream tasks, including Chronos-T5 Ansari et al.
+(2024) and Choronos-Bolt (https://huggingface.co/autogluon/chronos-bolt-base), PatchTST Nie et al.
 (2023), Moirai-1.0-R Woo et al.
-(2024), Chronos-T5 Ansari et al.
-(2024) and Choronos-Bolt (https://huggingface.co/autogluon/chronos-bolt-base).
-The models use different architectures, time series tokenization techniques and hyperparameters for numeric downstream tasks as summarized in Table 2.
+(2024) and Lag-Llama Rasul et al.
+(2024).
 Table 2: LLM models architectures, time series tokenization techniques and hyperparameter choices.
@@
 Lag-Llama Decoder Lag Feature $L=32$
+The considered models use different architectures, time series tokenization techniques and hyperparameters for numerical downstream tasks.
+For instance, Lag-Llma use decoder only transformer, PatchTST and Moirai-1.0-R use vanilla Transformer enoder, while Chronos-T5 and Choronos-Bolt use encoder-decoder transformer.
+Different baselines achieve contextual embedding in different ways.
+For example, PatchTST focuses on tokenizing time series as patches and uses self-attention for modeling dependencies within each patch and across patches, while CHORNOS-T5 and CHRONOS-Bolt adapt language modeling architectures minimally and generate categorical tokens by applying scaling and quantization.
+The details of these baselines are summarized in Table 2.
 Table 3: Real and Synthetic Datasets
@@
 We also illustrate our findings using KernelSynth Ansari et al.
-(2024) (see Algorithm 1 in Appendix E for details), a method that generates $10$ more synthetic datasets via Gaussian processes in Section 5.
-As in Section 4.1, we select two different datasets from each numerical domain (as shown in Table 3) and then perform qualitative analysis with synthetic datasets and quantitative analysis with real datasets.
-The results of both of this analysis are provided in the next two sections.
-Figure 6: NMSE vs isotropy analysis for $10$ different synthetic datasets of $5$ different domains.
+(2024) (see Algorithm 1 in Appendix E for details), a method that generates $10$ additional synthetic datasets via Gaussian processes in Section 5.
+We select two different datasets from each numerical domain (as shown in Table 3) and then perform qualitative analysis with synthetic datasets and quantitative analysis with real datasets.
+The results of these analyses are provided in the next two sections.
 5.1 Qualitative Analysis
-In this section, we analyze the time series forecasting by the baseline LLMs qualitatively.
+We now analyze the time series forecasting by the baseline LLMs qualitatively.
 We focus on synthetically generated time series for a controlled analysis of different types of time series patterns which belong to $5$ different domains, such as linear, seasonality, trend, non-linear and stochastic.
@@
 So all isotropic measure provided in this section is based on the last layer of the baselines.
+Figure 3: NMSE vs isotropy analysis for $10$ different synthetic datasets of $5$ different domains.
 We begin by analyzing time series forecasting performance (i.e., NMSE) for different baselines and its relation with isotropy in Figure 3.
 For instance, in Figure 3 b, we have (NMSE = $0.0000066$ and cosine similarity = $|-0.00076|$) for seasonality-Dataset 1 and (NMSE = $0.00012$ and cosine similarity = $0.0047$) for seasonality-Dataset 2 for Chronos-T5.
-This shows stronger isotropy exists (i.e., inter-type cosine similarity value is close to 00) in Chronos-T5’s embedding space for seasonality-Dataset 1 which preserves the structure in its hidden representations and causes good downstream task performance.
-On the other hand, a weaker isotropy exists (i.e., inter-type cosine similarity value is far from 00) in Chronos-T5’s embedding space for seasonality-Dataset 2, and thus causes a lack of structure in its hidden representations which leads to bad downstream task performance as compared to seasonality-Dataset 1.
-The NMSE and inter type cosine similarity may also vary for different language models and datasets.
-For example, in Figure 6 c, we can see that the the NMSE for trend-Dataset 1 is lower for PatchTST and Morai, while higher for Chronos-T5, Chronos-Bolt and Lag-Llma.
-In contrast, the NMSE for trend-Dataset 2 is lower for Chronos-T5, Chronos-Bolt, and Lag-Llma, while higher for PatchTST and Morai.
-A similar analysis can also be observed for other synthetic datasets and baselines in Figure 6 b, Figure 6 d, and Figure 6 e.
-This shows that any dataset from any particular domain may cause different forecasting performances for different baselines,
-Figure 7: NMSE vs isotropy comparison across different input context lengths for synthetic datasets.
-as it generates different hidden representations in contextual embedding spaces (see Appendix F for full visualization) for language models with different architectures and tokenization techniques.
+This shows that stronger isotropy exists (i.e., inter-token cosine similarity value is close to 00) in Chronos-T5’s embedding space for seasonality-Dataset 1 which preserves the structure in its hidden representations and causes good downstream task performance.
+On the other hand, a weaker isotropy exists (i.e., inter-token cosine similarity value is far from 00) in Chronos-T5’s embedding space for seasonality-Dataset 2, which, in turn, causes a lack of structure in its hidden representations, thereby leading to bad forecasting performance as compared to seasonality-Dataset 1.
+The NMSE and inter-type cosine similarity can also vary across different language models and datasets.
+For example, in Figure 3c, the NMSE for trend-Dataset 1 is lower for PatchTST and Moirai, but higher for Chronos-T5, Chronos-Bolt, and Lag-Llama, compared to their respective NMSE on trend-Dataset 2.
+Conversely, for trend-Dataset 2, the NMSE is lower for Chronos-T5, Chronos-Bolt, and Lag-Llama, but higher for PatchTST and Moirai, compared to their respective NMSE on trend-Dataset 1.
+A similar analysis can also be observed for other synthetic datasets and baselines in Figures 3 b, 3 d, and 3 e.
+Figure 4: Variations in Chronos-T5’s hidden representations for different input context lengths for the same synthetic dataset non-linear-Dataset 1 : (a) Contextual embedding space for input context length $L=500$.
+(b) Contextual embedding space for input context length $L=100$.
+This shows that any dataset from any particular domain may cause different forecasting performances for different baselines, as it generates different hidden representations (see Appendix F for full visualization) in contextual embedding spaces, and hence, different isotropy measures.
 Next, we examine the influence of isotropy on forecasting performance in two important scenarios: a) different input context lengths, and b) different levels of noises in the input data.
@@
 Isotropy in different input context lengths.
-We first analyze the impacts of isotropy under varying input context lengths.
+We first analyze the effect of isotropy under varying input context lengths.
+We begin with an illustration in Figure 4 where we show how the hidden representations of Chronos-T5 vary for two different input context lengths, such as $L=500$ and $L=100$, for non-linear-Dataset 1, which generates different isotropic measures for different input context lengths.
+Figure 5: NMSE vs isotropy comparison across different input context lengths for synthetic datasets.
 In Figure 5, we compare the NMSE vs isotropy across two different input context lengths, $L=500$ and $L=100$, for different synthetic datasets.
@@
 In this scenario, the isotropy increases for the input context length $L=100$, which causes the decrease in NMSE for chornos-T5.
-In Figure 11 in Appendix F, we show how the hidden representations of Chronos-T5 vary for two different input context lengths, such as $L=500$ and $L=500$, for non-linear-Dataset 1, which generates different isotropic measures for different input context lengths.
 In practice, the input context length is often selected randomly or through trial and error, which may cause higher forecasting errors for different datasets.
@@
 Again, we use the Chornos-T5 as an example language model.
-In Figure 8, we compare the NMSE vs isotropy across two different noise levels, one without noise, and the other with Gaussian noise with $\sigma=0.05$ standard deviation.
-From Figure 8, we can see consistently lower isotropy (i.e., inter-type cosine similarity far from 00) for all noisy synthetic datasets as compared to the datasets without noise.
+Figure 6 compares the NMSE vs isotropy across two different cases, one without noise, and the other with Gaussian noise with a standard deviation $\sigma=0.05$ standard deviation.
+From Figure 6, we can see consistently lower isotropy (i.e., inter-token cosine similarity far from 00) for all noisy synthetic datasets as compared to the datasets without noise.
 For instance, in trend-Dataset 2, we have (NMSE$=0.000024$, cosine similarity$=|-0.00022|$) and (NMSE$=0.0012$, cosine similarity$=0.0040$) for $\sigma=0$ and $\sigma=0.05$, respectively.
@@
 In practice, many real-world numerical domains—such as those in nature and energy—exhibit noisy and dynamic behavior.
+Figure 7: NMSE vs isotropy analysis for $12$ different real datasets of $6$ different domains.
 In these environments, it is often infeasible to measure noise in real time or to pre-process the input time series for improved performance.
-However, the isotropy in LLMs hidden representations can be readily measured, and thus, can be leveraged to enhance forecasting performance by identifying and mitigating the effects of noisy inputs in contextual embedding space.
+However, the isotropy in the hidden representations of LLMs can be readily measured, and thus, can be leveraged to enhance forecasting performance by identifying and mitigating the effects of noisy inputs in contextual embedding space.
 5.2 Quantitative Analysis
-In this section, we present our main results on $12$ real datasets which belong to $6$ different numerical domains such as energy, nature, finance, healthcare, retail and transportation.
+Next, we present our main results on $12$ real datasets which belong to $6$ different numerical domains including energy, nature, finance, healthcare, retail, and transportation.
 As our qualitative analysis in Section 5.1, we select two different datasets from each numerical domain and the isotropy measure from LLM’s last layer to show the impact of isotropy on NMSE performance for different language models.
-Figure 9: NMSE vs isotropy analysis for $12$ different real datasets of $6$ different domains.
 In Figure 7, we analyze the time series forecasting performance of different baselines and its relation with isotropy for different real datasets.
-For instance, in Figure 9 e, we have (NMSE = $0.0061$ and cosine similarity = $0.0020$) for retail-Dataset 1 and (NMSE = $0.1255$ and cosine similarity = $0.1931$) for retail-Dataset 2 for Morai.
-This indicates the existence of stronger isotropy in Morai’s embedding space for retail-Dataset 1 which preserves the structure in its hidden representations and causes good downstream task performance.
-On the other hand, a weaker isotropy exists in Morai’s embedding space for retail-Dataset 2, and thus causes a lack of structure in its hidden representations which leads to bad downstream task performance as compared to retail-Dataset 1.
-The NMSE and inter type cosine similarity may also vary for different real datasets and language models.
-For example, in Figure 9 a, we can see that the NMSE for energy-Dataset 1 is lower for Chronos-T5, Chronos-Bolt, PatchTST and Morai, while higher Lag-Llma.
-In contrast, the NMSE for energy-Dataset 2 is lower for Morai while higher for other baselines.
-A similar analysis can also be observed for other synthetic datasets and baselines in Figure 9 b and Figure 9 c.
-This again shows that datasets from the same numerical domain can cause varying forecasting performance across different baselines, as they generate distinct hidden representations in contextual embedding spaces depending on the language model architecture and tokenization strategy.
+For instance, in Figure 7 e, we have (NMSE = $0.0061$ and cosine similarity = $0.0020$) for retail-Dataset 1 and (NMSE = $0.1255$ and cosine similarity = $0.1931$) for retail-Dataset 2 for Moirai.
+This indicates the existence of stronger isotropy in Moirai’s embedding space for retail-Dataset 1 which preserves the structure in its hidden representations and causes good downstream task performance.
+On the other hand, a weaker isotropy exists in Moirai’s embedding space for retail-Dataset 2, which yields a lack of structure in its hidden representations and, consequently, bad downstream task performance as compared to retail-Dataset 1.
+The NMSE and inter-type cosine similarity can vary across different real datasets and language models.
+For example, in Figure 7a, the NMSE for energy-Dataset 1 is lower for Chronos-T5, Chronos-Bolt, PatchTST, and Moirai, but higher for Lag-Llama, compared to their respective NMSE on energy-Dataset 2.
+Conversely, the NMSE for energy-Dataset 2 is
 Figure 8: NMSE vs isotropy comparison across different input context lengths for real datasets.
+lower for Moirai but higher for the other baselines, compared to their respective NMSE on energy-Dataset 1.
+A similar analysis can also be observed for other synthetic datasets and baselines in Figure 7 c and 7 e.
+This again shows that datasets from the same numerical domain can cause varying forecasting performance across different baselines, as they generate distinct hidden representations in contextual embedding spaces, and hence, different isotropy measures, depending on the language model architecture and tokenization strategy.
 Finally, in Figure 8, we compare the NMSE vs isotropy for varying input context lengths to observe its impact on the real datasets.
@@
 We compare the results for two different input context lengths: 1) the recommended input context length $L=144$ and the reduced input context length $L=96$.
-As can be seen from the figure, the inter-type cosine similarity values become far from 00, i.e., from $0.0097$ to $0.0220$ for energy-Dataset 1 and from $0.0026$ to $0.0103$ for transport-Dataset 1, which in turn decreases the NMSE performances.
-On the other hand, the inter-type cosine similarity values become close to 00, i.e., from $0.1091$ to $0.0112$ for nature-Dataset 2 and from $0.2014$ to $0.0133$ for finance-Dataset 1, which in turn improves the NMSE performances.
+As can be seen from the figure, the inter-token cosine similarity values become far from 00, i.e., from $0.0097$ to $0.0220$ for energy-Dataset 1 and from $0.0026$ to $0.0103$ for transport-Dataset 1, which in turn decreases the NMSE performances.
+On the other hand, the inter-token cosine similarity values become close to 00, i.e., from $0.1091$ to $0.0112$ for nature-Dataset 2 and from $0.2014$ to $0.0133$ for finance-Dataset 1, which in turn improves the NMSE performances.
 Thus, the variation in the recommended input context length may not only decrease the NMSE performances, but can also increase for some datasets.
 6 Conclusion and Limitations
-Isotropy in embeddings as studied here can serve as a foundation for future research on the deeper understanding of LLMs and their applications in various domains.
-Beyond isotropy, there could be other methods to approximate the partition function with a constant and make the logits useful for the numeric downstream tasks.
-Moreover, our isotropy study only ensured the existence of structure in the LLM hidden representations and provides a performance guarantee when the structure is preserved by isotropy.
-Improving the numeric downstream task performance when structure is not preserved in the LLM representations is a topic of future work.
-Appendix A Proof of Theorem 3.1
-Theorem A.1.
+In this work, we introduced a novel approach to investigate the role of isotropy in LLM hidden representations for numerical downstream tasks.
+By deriving an upper bound for the Jacobian matrix which collects all first-order partial derivatives of self-attention with respect to the input pattern, we showed that the self-attention mechanism implicitly aligns with the dominant eigenvectors of the input correlation structure and induces isotropy in the contextual embedding space.
+The existence of isotropy in the contextual embedding space was found to stabilize the partition function and enable better generalization in numerical downstream tasks across different models and datasets.
+Our empirical analysis across $10$ synthetic and $12$ real numerical datasets, and $5$ different language models further validated the consistent relationship between isotropy and forecasting performance, highlighting isotropy as a reliable indicator of structured representation learning.
+These insights open up a new interpretability frontier for LLMs in numerical domains.
+While isotropy offers a principled way to preserve useful structure, there may be alternative approaches to approximating the partition function and guiding numerical reasoning.
+Moreover, developing mechanisms to recover or enhance structure when isotropy is weak remains an important avenue for future work.
+Ultimately, we believe that incorporating structural insights like isotropy into the LLM design pipeline can significantly improve their reliability and adaptability to numerical domains.
+Appendix A Proof of Theorem 1
+Theorem 1.
 Let the logits of the ground-truth model be bounded.
 Then for any $f^{*}(k,l)$, there exists a set of functions $\{\hat{z}_{i}(k,l)\}_{i=1}^{|\mathcal{V}|}$ such that for all $k$ and $T_{l+1}$, the predictive distribution of the student model $\hat{p}\,(k_{T_{l+1}}\,|\,\mathbb{k}_{1:T_{l}})$ matches that of ground-truth model $p^{*}(k_{T_{l+1}}\,|\,\mathbb{k}_{1:T_{l}})$ and $\hat{f}(k,l)=0$.
-In other words, there exists a student model with the same pre-training loss as the ground-truth model, but its logits are ineffective for the numeric downstream tasks.
+In other words, there exists a student model with the same pre-training loss as the ground-truth model, but its logits are ineffective for the numerical downstream tasks.
 Proof.
@@
 this implies that $\sigma(\hat{z}_{j}(k,l)-b_{j}^{*})=0$. Hence, $\forall k,T_{l+1}$ and we have $\hat{f}(k,l)=0$. ∎
-Appendix B Proof of Lemma 4.1
-Lemma B.1.
+Appendix B Proof of Lemma 1
+Lemma 1.
 Consider the Jacobian matrix $\mathbb{J}=\left[\frac{\partial g_{i}(\mathbb{\Psi})}{\partial\mathbb{\psi}_{j}}\right]_{i,j=1}^{|\mathcal{V}|}$, which represents the gradient of the self-attention mapping $G(\mathbb{\Psi})$ with respect to the input time series token embeddings.
@@
 $=|\mathbb{\Lambda}|_{2}\sum_{i=1}^{|\mathcal{V}|}\left(p_{i,i}+\frac{1}{2}\right)|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}p_{i,:}|^{2}+|\mathcal{V}|+|\mathbb{\Lambda}|_{2}\sum_{i\neq j}^{|\mathcal{V}|}p_{i,j}\left|\mathbb{\psi}_{j}-\sum_{q=1}^{|\mathcal{V}|}p_{i,q}\mathbb{\psi}_{q}\right|^{2}+\frac{|\mathbb{\Lambda}|_{2}}{2}\sum_{j=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}|^{2}$
-The theorem below shows that $\mathbb{\Lambda}$ minimizing the objective $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}|^{2}$ contains the largest $m$ eigenvectors of the correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$ of input time series token embeddings where $m$ is the rank of $\mathbb{\Lambda}$.
+Theorem 2 shows that $\mathbb{\Lambda}$ minimizing the objective $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}|^{2}$ contains the largest $m$ eigenvectors of the correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$ of input time series token embeddings where $m$ is the rank of $\mathbb{\Lambda}$.
 Lemma 1 implies that one of the key components in the Jacobian’s upper bound takes the form $|\mathbb{\psi}_{i}-\sum_{j=1}^{|\mathcal{V}|}p_{i,j}\mathbb{\psi}_{j}|^{2}$.
@@
 Assuming all input time series patterns are zero-centered, i.e., $\bar{\mathbb{\psi}}=0$, we have $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}|^{2}=\text{tr}\left((I-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda})^{2}\mathbb{\Psi}^{\top}\mathbb{\Psi}\right)$.
-Theorem 4.1 establishes that the optimal $\mathbb{\Lambda}$ that minimizes $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}|^{2}$ is spanned by the top $m$ eigenvectors of $\mathbb{\Psi}^{\top}\mathbb{\Psi}$, where $m$ equals the rank of $\mathbb{\Lambda}$.
-Appendix C Proof of Theorem 4.1
-Theorem C.1.
+Theorem 2 establishes that the optimal $\mathbb{\Lambda}$ that minimizes $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}|^{2}$ is spanned by the top $m$ eigenvectors of $\mathbb{\Psi}^{\top}\mathbb{\Psi}$, where $m$ equals the rank of $\mathbb{\Lambda}$.
+Appendix C Proof of Theorem 2
+Theorem 2.
 Let the eigenvalues of the correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$ be ordered as $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{D}$, and let $\gamma_{i}\in\mathbb{R}^{D}$ for $i=1,\ldots,D$ denote their associated eigenvectors.
@@
 Appendix F Full Visualization of PCA plots for different models
-The Variations in Chornos-T5’s hidden representations for different input context lengths is depicted in Figure 11
-Figure 11: Variations in Chronos-T5’s hidden representations for different input context lengths for the same synthetic dataset non-linear-Dataset 1 : (a) Contextual embedding space for input context length $L=500$.
-(b) Contextual embedding space for input context length $L=100$.
 The full visualization of PCA plots of different models is provided below. We use the synthetic Dataset 1, and Dataset 2 from non-linear domain for illustration.